### v1

# A General Class of Model-Free Dense Precision Matrix Estimators

Abstract
We introduce prototype consistent model-free, dense precision matrix estimators that have broad application in economics. Using quadratic form concentration inequalities and novel algebraic characterizations of confounding dimension reductions, we are able to: (i) obtain non-asymptotic bounds for precision matrix estimation errors and also (ii) consistency in high dimensions; (iii) uncover the existence of an intrinsic signal-to-noise—underlying dimensions tradeoff; and (iv) avoid exact population sparsity assumptions. In addition to its desirable theoretical properties, a thorough empirical study of the S&P 500 index shows that a tuning parameter-free special case of our general estimator exhibits a doubly ascending Sharpe Ratio pattern, thereby establishing a link with the famous double descent phenomenon dominantly present in recent statistical and machine learning literature.
Keywords: Precision Matrix Estimation; Double Descent; Latent Space; Hidden Factors.
JEL Code: C18, C21, C55.
1 Introduction
In this paper, we propose a general class of new estimators for precision matrices in settings with a large—and potentially diverging—number of observed variables that is dense and model-free. Hence, we neither use penalized estimation nor estimate an imposed model to generate a precision matrix estimator.
Our main contributions are as follows. We introduce a general prototype class of consistent-non-sparse and unpenalized precision matrix estimators based on a deep connection between matrix algebra and hidden factor analysis. Hidden factor analysis is used only for achieving consistency in high dimensions via dimension reduction. Our estimators are model-free, and do not need to estimate hidden factors. Moreover, we derive non-asymptotic upper bounds for estimation errors that allow us to see the effects of a new complexity random variable, number of confounding factors, and signal-to-noise ratio. Our precision matrix estimator has three broad areas of interest. First, it can be used to construct debiased estimators for parameters in high-dimensional regression models. Debiased estimators are essential for constructing valid confidence intervals in high-dimensional settings (see Caner and Kock (2018), van de Geer et al. (2014)). A key challenge in this context is obtaining an estimate of the inverse of the Gram matrix, as the sample second moment matrix is typically singular when the number of variables exceeds the sample size. Our estimator provides a viable alternative for estimating this inverse, and can be readily adapted for use in such models. Second, in network analysis, under normally distributed random outcome variables, conditional dependence between two random variables is represented as a nonzero term in the precision matrix of all outcome variables. Our method avoids imposing “artificial” conditional independence assumptions—i.e., forcing zero entries in the precision matrix that correspond to the absence of links between variables. Instead, it preserves all connections in the network, allowing some to be minor and others to be dominant. A direct application of such setting is the study of tax interactions between US states in the study of De Paula et al. (2024). They impose sparsity of their social interaction matrix i.e., zero interactions between states (see equations (8) and (10) of their paper). Third, in financial econometrics, our estimator can be used where precision matrices of asset returns play a central role in tasks such as optimal portfolio allocation, portfolio variance estimation, and Sharpe Ratio calculation. A detailed literature review and discussion of these three broad applications is provided in Section 4.4.
1.1 Brief Overview of Estimators
We start by providing a formula for population precision matrix that can tie each variable to the other variables in a system. Then, to bypass the high-dimensionality of the problem, our conceptual innovation is to express each variable as a linear function of common but hidden factors, and solve for the corresponding precision matrix estimator. In this sense, hidden factors act as confounders. We make no identifying assumptions about hidden factors, nor estimate them.
As special subcases of our general estimator, we obtain a new adaptive principal components (adaptive PCR from now on) estimator and the ridgeless regression estimator (RRE from now on). We prove consistency of both estimators when the number of outcomes exceeds the sample size. Empirically, we find that the Sharpe Ratio of portfolios constructed using the ridge-less regression estimator exhibits a double ascent pattern in the S&P 500. This is significant, as it demonstrates that profitable portfolios can be formed without the need to estimate any tuning parameters.
We compare two subcase estimators: the ridge-less regression estimator (RRE) and adaptive principal component regression (PCR). In finite samples, the RRE has one fewer bias term than adaptive PCR, giving it a potential advantage. However, asymptotically the ranking may reverse, since the bias in adaptive PCR depends on the number of factors, while the bias in RRE depends on sample size. For instance, because the number of factors is typically smaller than the sample size, adaptive PCR may outperform RRE in estimating individual rows of the precision matrix.
We support these insights through simulations and empirical analysis. Specifically, we find that our ridge-less precision matrix estimator achieves a Sharpe Ratio of 0.2203 for a large portfolio of 357 S&P 500 stocks during the out-of-sample period from January 2015 to December 2023, after accounting for transaction costs. The closest competitor is the nodewise regression-based approach of Callot et al. (2021), which attains a Sharpe Ratio of 0.2082.
1.2 Dimension Reduction
Why do we use a hidden factor setup? First of all, we show that precision matrix estimation is related to linear regression and that can be related to hidden factors (see Online Appendix B). Each economic or financial variable can be related to latent factors and this can be thought as confounding variables as in causality related literature such as in Kaltenpoth and Vreeke (2023), Angrist and Pischke (2009), Imbens and Rubin (2015). Also the true variables of interest may be measured with errors, and latent factors can represent the ”true” variables rather than the reported ones, especially in emerging markets. Variable such as culture, sentiment, market can be extremely valuable in certain macro, financial data sets. These hidden factors in a linear model can be used to explain international stock returns such as in Singapore, Malaysia, USA as seen in Brooks and Del Negro (2004).
Another key aspect of using hidden factors is their effectiveness as a tool for dimension reduction in high-dimensional econometrics. This stems from the fact that the number of observed outcomes typically exceeds the number of underlying hidden factors. In our paper, we allow both the number of outcomes and the number of hidden factors to grow with the sample size. The role and significance of latent factors and the estimation of their key components, factors and loadings, have been largely extended and further linked to asset pricing in Lettau and Pelger (2020), causal inference and estimation of counterfactuals or treatment effects in Athey et al. (2019); Athey et al. (2021); Yan and Wainwright (2024); Kallus et al. (2018), recommenders systems in Agarwal et al. (2023), macroeconomic forecasting in Stock and Watson (2002a, b, 2016), matrix completion and missing panel data in Xiong and Pelger (2023); Cahan et al. (2023); Bai and Ng (2021); Duan et al. (2024); Bryzgalova et al. (2023). Chen et al. (2021) discuss how latent factors can model important features of network data such as reciprocity, degree heterogeneity, and homophily. They apply these methods to the estimation of gravity equations of trade between countries, showing how unobserved latent factors significantly influence observed trade volumes, even after controlling for traditional gravity variables like distance. Excellent theoretical and practical algorithmic results obtained in all these works usually serve as a strong indication that there is indeed a very strong rationale for utilization of latent factors. Somewhat paradoxically, while we will rely on similar factor modeling concepts and underlying intuitions that they bring along, we will be able to design new and novel precision matrix estimators without actually focusing on directly estimating factors themselves.
1.3 A Brief Discussion of Sparsity in Economics
Our proposed class of estimators is broadly applicable across a range of settings, including those where the underlying data do not exhibit sparsity. For instance, Giannone et al. (2021) analyze six widely used datasets in macroeconomics, microeconomics, and finance, and find that dense models consistently outperform sparse ones in terms of out-of-sample predictive accuracy. These applications include: (i) forecasting U.S. economic activity, (ii) identifying determinants of economic growth across countries, (iii) explaining the decline in crime rates across U.S. states, (iv) analyzing judicial rulings on eminent domain across U.S. circuits, (v) predicting the U.S. equity premium, and (vi) accounting for the cross-sectional variation in U.S. stock returns.
In asset pricing, the assumption of sparsity for explaining or predicting economic variables is particularly strong and has been critically examined (see p.33 and Section 3.7 of Nagel (2021)). In that study, sparsity is imposed between the outcome variable and the predictors, effectively forcing some coefficients to be zero—even if those predictors have a non-negligible but small relationship with the outcome. This can be problematic in settings where many variables contribute modestly to predictive performance. Techniques like Lasso may shrink all small coefficients to zero, resulting in an artificially sparse model and a significant loss of predictive power. [^1] Figure 3.2 of Nagel (2021) illustrates how sparsity-based methods can underperform due to this limitation.
1.4 Literature Survey
Several studies in the statistics literature have studied precision matrix estimation. An excellent survey is by Cai et al. (2016). In Section 3 of their paper, they provide two major approaches to precision matrix estimation. The first approach is neighborhood-based and closely related to nodewise regression, where a lasso regression is performed for each variable against all others. Alternatives to lasso, such as the Dantzig selector, can also be used within this framework (see Cai and Liu (2011), Callot et al. (2021)). The second major approach is based on penalized likelihood estimation. As shown by Bertsimas et al. (2020), lasso-type penalties in this context can induce shrinkage bias and lead to overly sparse precision matrix estimates, with many zero entries in each row. In a penalized likelihood based setting, they reformulate the problem as a mixed-integer optimization problem, which is solvable using convex optimization techniques. This approach produces better results than lasso based approaches. Sparsity of precision matrix is used in these two main approaches.
A distinct approach, commonly used in finance and statistics, derives the precision matrix by inverting the covariance matrix using the Sherman–Morrison–Woodbury (SMW) formula. In this framework, sparsity is assumed in the covariance matrix of the idiosyncratic errors, while asset returns are modeled through a factor structure (see Fan et al. (2013), Caner et al. (2023)).
Our approach is different from the three approaches discussed above, and provides a new way of solving the same problem. Compared to the first two, we are not assuming exact sparsity in the precision matrix of outcomes. Compared to the last one, we are not using SMW formula to invert the covariance matrix. This allows us to bypass imposing a structure on the outcomes utilized in the covariance inversion. Instead we use a new precision matrix formula without sparsity and rely on hidden factors to generate each outcome. This effectively ensures that we do not benefit from the SMW matrix inversion formula. To be clear, we use model-free dense estimation, but consistency in high dimensions is achieved through hidden factor based dimension reduction without a need to estimate factors or identify them.
Note that nodewise regression based estimator of Callot et al. (2021) uses sparsity of precision matrix of outcomes. Then residual nodewise regression estimator of Caner et al. (2023) crucially depends on observed factors, and imposes sparse precision matrix of errors. Technique used in those papers cannot extend to hidden factors since observed factor structure in Caner et al. (2023) uses OLS estimation which is not applicable to more complicated estimation problems such as hidden factor structure. Note that Callot et al. (2021) has no factor structure and uses lasso based regression to estimate the precision matrix of outcomes.
We also tie our results to signal-to-noise ratio aspect which is not available in Callot et al. (2021), Caner et al. (2023) papers. We show that to get consistency of our estimators, signal-to-noise ratio should be larger than number of factors. Our method can be seen as a complementary approach to nodewise regression of Meinshausen and Bühlmann (2006). Nodewise regression uses an $l_{1}$ norm based penalization to form the rows of the precision matrix. Doing so, the underlying assumption is the imposition of sparsity in the precision matrix which may be interpreted as a strong assumption.
An excellent paper by Fan et al. (2013) provides a statistical high-dimensional analysis of covariance matrix of outcomes via hidden factors based estimation. Similar to us, it also relies on hidden factors but uses a thresholding based principal components estimator for the covariance matrix, and then inverts this through SMW formula. The key difference is that our estimator is model-free and uses hidden factors for consistency proof only. There are also other technical and conceptual differences that are worth highlighting. First, our proposed class of estimators is general and in addition to the PCR based ones it also includes the ridge-less regression estimator (RRE). Second, our signal assumption is much weaker as we do not need the minimum signal to grow at the same rate as the number of assets in the portfolio. Also, we allow for diverging noise assumption (this is a notable contrast to Fan et al. (2013) where the noise is tied to the maximum eigenvalue of the covariance matrix of errors with assumed constant dimension $p$), which is reflective of financial markets. Moreover, as the root of our approach resembles a finite sample analysis, it also allows to explicitly tie all the estimation errors to signal-to-noise ratio thereby making its importance in the rate of convergence estimates clearly visible. In particular, we uncover that the signal-to-noise ratio that dominates the number of factors ensures consistency of our precision matrix estimator.
Our paper differs from Bing et al. (2021) and Bunea et al. (2022) in a fundamental way. These papers analyze general linear class of estimators in hidden factor models in linear regression model and is only interested in risk of the estimators. Their papers consider only one outcome variable and how it can be predicted with this new general class of estimators. Theorem 3 of Bing et al. (2021) specifically provides a risk upper bound and this bound depends on the magnitude of coefficients in a linear system. Our analysis on the other hand, considers the precision matrix estimation of outcome variable in linear regression via hidden factors (i.e. an application is large portfolio formation). We show how matrix algebra can be combined with latent factors to provide a new precision matrix formula in order to achieve consistency. Then a novel idea of estimation is to use a linear linkage between outcome variables to obtain precision matrix estimator based on the matrix algebraic new formula. One of the main difficulties, theoretically, is that the estimation of the main diagonal terms in the precision matrix estimation. The proof technique and ideas in Bing et al. (2021), Bunea et al. (2022) do not simply extend to this case. Hence, the precision matrix estimation is different than the risk analysis of a single outcome-asset. We also provide an explicit connection in our estimation error upper bounds to signal-to-noise ratio. The results in subcases are different also in a risk analysis versus precision matrix estimation. Bing et al. (2021), Bunea et al. (2022) show that in certain very high dimensional cases, ridge-less regression (RRE) can perform better than adaptive PCR in terms of risk asymptotically. However, in case of estimation errors of rows of precision matrix errors (in $l_{2}$ norm) we show adaptive PCR can perform better than RRE even in very high dimensional cases. These are due to different proofs for risk versus precision matrix estimation.
1.5 Notation
We consider notation now. Define $C,c$ as generic positive constants. Note that transpose of any matrix $A:d\times l$ is represented by $A^{\prime}$, and transpose of a column vector, $a$ is represented by $a^{\prime}$ unless it is stated otherwise for notational convenience. Let $\|A\|_{2}$ represent the spectral norm of matrix $A$, and $\|a\|_{2}$ the Euclidean norm of a vector $a$. Also, we denote by $\|A\|_{F}$ the Frobenius norm of $A$. Let $\|A\|_{l_{\infty}}:=\max_{1\leq j\leq d}\|A_{j}\|_{1}$ be the maximum row sum norm, with $A_{j}:l\times 1$ as the $j$ th row among $d$ rows in matrix A, written in column format. The row format for $j$ th row of $A$ is $A_{j}^{\prime}:1\times l$. Let $\|A\|_{\infty}:=\max_{i,j}|A_{ij}|$, with $A_{ij}$ representing $i,j$ cell of matrix A. Denote the $k$ th largest singular value of $A$ matrix as $\sigma_{k}(A)$ with $k=1$ being the largest singular value. Let $A^{+}$ represent the Moore-Penrose inverse of matrix A. Let $\lambda_{k}(M)$ represent the $k$ th largest eigenvalue of $M:d\times d$, a square matrix, except $k=1,k=d$. However, for ease of understanding, the largest eigenvalue of $M$ is denoted as $Eigmax(M)$ and the minimum eigenvalue of $M$ is $Eigmin(M)$. $tr(M)$ represents the trace of matrix $M$. Let $a_{n}\asymp b_{n}$ show the same order for $a_{n},b_{n}$ random variables, i.e. $a_{n}=O_{p}(b_{n})$, and $b_{n}=O_{p}(a_{n})$. Section 2 proceeds with the idea of dimension reduction through hidden factors to achieve high dimensional consistency in the proofs. Section 3 introduces the estimators. Section 4 has consistency results. Section 5 provides empirical evidence. Conclusion shows future possible extensions of our work. Appendix shows the proof of Theorem 2. Simulations, some of the empirics and rest of the proofs are relegated to Online Appendix.
2 Dimension Reduction
In our setup we have $p$ outcome variables, and they are explained by $K$ common hidden factors to ensure dimension reduction, with $K<<p$. This will help in asymptotics and achieving consistency, however our estimators will be model-free and will not use hidden factors estimators. Our estimators will only use outcome variables and this will be clear in the next section. Specifically, our estimators are defined in (11)(12). Let $Y$ represent the $n\times p$ matrix of outcomes, each row representing the $t$ th time period, $t=1,\cdots,n$, and $j=1,\cdots,p$ representing the outcomes. $F$ represents $n\times K$ matrix of hidden factors, where each row, $t=1,\cdots,n$ represents a time period, and columns represent the hidden factors, $k=1,\cdots,K$. $A$ represents $K\times p$ matrix of factor loadings, and $U:n\times p$ as the error matrix. We have the following compact form of the linear hidden factor structure for dimension reduction (which will be decomposed via rows of Y into two equations immediately below in (2)-(3)).
$Y=FA+U.$ (1)
In accordance with our analysis, we subdivide all $p$ outcomes into $j$ th one and all the remaining ones. In that respect, start by defining each row of the outcome as $y_{t}:p\times 1$ (represents each row, in column form, whereas $y_{t}^{\prime}:1\times p$ is the row vector form) of $Y$ above, and $f_{t}:K\times 1$ represents each row (in column form) of $F$, and $u_{t}:p\times 1$ is the $t$ th row (in column form) of U. We decompose $A$ into $a_{j}:K\times 1$ factor loading for outcome j, and $A_{-j}:p-1\times K$ as factor loading matrix for all other outcomes except $j$. In other words $A:=[a_{j},A_{-j}^{\prime}]$.
We model $y_{jt}$ as the $j$ th outcome variable at time period $t=1,\cdots,n$. As an example, the $j$ th outcome here can be $j$ th asset’s excess return (over risk free rate such as 3 month T-Bill in USA) at time $t$. We define all the outcomes except $j$ th outcome as $Y_{-jt}:p-1\times 1$ vector at time $t$.
We decompose (1) into $j$ th outcome and the rest. Keeping all of the above in mind, an interesting parallel with the hidden (latent) factor regression in Bunea et al. (2022) and Bing et al. (2021) can be made as well.222 There in (1) of Bing et al. (2021) they use regressors to predict the outcome, and all of the regressors and outcome is a linear function of hidden variables. In the $t$ th row of $Y$ in (1), $j$ th outcome can be written as
$y_{jt}=f_{t}^{\prime}a_{j}+u_{jt},$ (2)
and we write all the outcomes except from the $j$ th one, in $t$ th row as
$Y_{-jt}=A_{-j}f_{t}+U_{-jt}.$ (3)
We use (2)(3) in our proofs rather than (1), and in this way our proofs are clear and tractable. Common hidden factors affect the $j$ th outcome in (2) as well as all remaining outcome variables in (3). Let $u_{jt},U_{-jt}$ represent the zero-mean errors for the $j$ th outcome and all other remaining outcomes, respectively.
In vector form, the j-th column of $Y$ in (1) can be written as
$y_{j}=Fa_{j}+u_{j},$ (4)
where $y_{j}:n\times 1$, $F:n\times K$ with each row as $f_{t}^{\prime}$. Then the remainder, except for the $j$ th column, of Y is:
$Y_{-j}=FA_{-j}^{\prime}+U_{-j},$ (5)
where $Y_{-j}:n\times p-1$ matrix, and $U_{-j}:n\times p-1$ matrix too. As we will see later on, the vector form representation given in (4)(5) will turn out to be beneficial in our proofs as well. Our assumptions are stated next. Before doing so, we briefly introduce a few additional notational conventions.
Define the variance of $j$ th error as $var(u_{jt})=\sigma_{j}^{2}\geq c>0$,where $c$ is a positive constant at time $t$, for $j=1\cdots,p$. Define the variance matrix of errors except the $j$ th one as $var(U_{-jt})=\Sigma_{U,-j}:p-1\times p-1$, for $j=1,\cdots,p$ at time $t$. So if $j=3$ for example $\sigma_{3}^{2}$ is the third asset error variance and $\Sigma_{U,-3}$ represents the covariance matrix of errors for all assets except the third one at time $t$. Define the covariance matrix of factors as $\Sigma_{f}:=cov(f_{t}):K\times K$. Define $y_{t}=(y_{jt},Y_{-jt}^{\prime})^{\prime}$ as the re-ordered $t$ th row of $Y$, and $\mu_{j}:=Ey_{jt}:1\times 1$, $\mu_{-j}:=EY_{-jt}:p-1\times 1$, and $\mu:=Ey_{t}:p\times 1$ at time $t$. Let the covariance matrix of outcomes be $\Sigma:=E(y_{t}-\mu)(y_{t}-\mu)^{\prime}$. Assumption 1 below will show iid nature of our data across $t=1,\cdots,n$. Define the precision matrix as $\Theta:=\Sigma^{-1}$. We provide Assumptions next.
Assumption 1.
(i). $y_{jt},Y_{-jt}$ are iid across $t=1,\cdots,n$ for each $j=1,\cdots,p$.
(ii). $u_{jt},U_{-jt}$ are zero mean random variable and vector respectively, which are iid across $t=1,\cdots,n$, for each $j=1,\cdots,p$.
(iii). $f_{t}$ are iid across $t=1,\cdots,n$.
(iv). $f_{t},u_{jt},U_{-jt}$ are mutually independent across $t=1,\cdots,n$ given $j=1,\cdots,p$.
Assumption 1 is a standard convention used in hidden factor literature (see, e.g., Bing et al. (2021) and Bunea et al. (2022).  [^3] Note also that (i)-(iii) implies that rows of F, Y, and U are iid. Note also, that independence of errors implies $Eu_{jt}U_{-jt}=0$ and $Eu_{t}u_{t}^{\prime}=I_{p},u_{t}=(u_{jt},U_{-jt}^{\prime})^{\prime}$ which is a sparsity restriction on the covariance matrix of errors. This is another assumption often present in the literature (see, e.g., Fan et al. (2008)) and amounts to a diagonal error covariance matrix. Relaxing this part of the assumption is certainly an interesting topic well worth of further exploring. However, since we pursue a general precision matrix of outcome variables in a hidden factor setup, and our outcomes are still correlated through the hidden factor structure, our extensive proofs would get more cumbersome. We therefore leave further work in this direction for a mathematically more narrowly focused future extension.
Note that this sparsity for the error covariance matrix does not imply sparsity of the covariance matrix of outcome variables in our framework. We are not imposing sparsity on the links between outcome variables as done in nodewise regression context.
Assumption 2.
(i). $\min_{1\leq j\leq p}Eigmin(\Sigma_{U,-j})\geq c>0$, where $Eigmin(\Sigma_{U,-j})$ represents the minimum eigenvalue of the matrix $\Sigma_{U,-j}$. Also $\max_{1\leq j\leq p}\sigma_{j}^{2}=\sigma^{2}\leq C<\infty$, and also $\min_{1\leq j\leq p}\sigma_{j}^{2}\geq c>0$.
(ii). Let $\delta_{n}:=\min_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}$, and $\delta_{n}$ is nondecreasing in $n$, and also $r_{n}:=\max_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}$, where $r_{n}$ is nondecreasing in $n$ but $r_{n}/p\to 0$, and clearly $\delta_{n}\leq r_{n}$.
Assumption 2(i) is also standard in high dimensional statistics literature. It allows for non-singular noise matrices, large noise magnitudes and divergence in the maximum eigenvalue of the (covariance of the) noise. This assumption is very much reflective of financial markets and is used in Caner et al. (2023).
Assumption 3.
(i). Define a positive constant $C_{p}>0$. Then assume the following relation between the number of outcomes, $p$ and the sample size $n$:
$p\leq n^{C_{p}/2}.$
(ii). For each $j=1,\cdots,p$, $A_{-j}:p-1\times K$ has full rank $K$, and $p>K+1$. We impose $\max_{1\leq j\leq p}\|A_{-j}\|_{2}=O(d_{1n})$ with $d_{1n}\to\infty$ as $n\to\infty$, and $d_{1n}/\sqrt{p}\to 0$. Also, $\Sigma_{f}$ has full rank $K$, and $Eigmax(\Sigma_{f})\leq C<\infty$, for C a positive constant.
With $C_{p}>2$, we can obtain $p>n$ and we cover high dimensional cases. In case of $0<C_{p}\leq 2$, we cover low dimensional cases, $p\leq n$. This assumption also allows $p=\gamma n$, with $\gamma\in[0,\infty)$. Full rank $\Sigma_{f}:=E(f_{t}-Ef_{t})(f_{t}-Ef_{t})^{\prime}:K\times K$ matrix is used in Bing et al. (2021), Bunea et al. (2022), reflecting $K<p-1$, where we assume the number of hidden factors to be less than number of assets minus one. We allow the spectral norm of $A_{-j}:p-1\times K$ matrix to be diverging in $n$, reflective of the high dimensional nature of $A_{-j}$, where both dimensions are allowed to grow.
Assumption 4.
(i). For each $j=1,\cdots,p$ let $u_{jt}$ be zero mean $\sigma_{j}\gamma_{e}$ subgaussian random variable which means for each $j=1,\cdots,p$ $Eexp(t_{1}u_{jt})\leq exp(t_{1}^{2}\sigma_{j}^{2}\gamma_{e}^{2}/2)$ for all $t_{1}\in R$. Define the zero mean vector $U_{-j,t}:=\Sigma_{U,-j}^{1/2}\tilde{U}_{-j,t}$, where $\tilde{U}_{-j,t}$ ($p-1\times 1$ vector) is $\gamma_{wj}$ sub-Gaussian vector which means for any unit vector ($p-1$ dimension) $v$, $v^{\prime}\tilde{U}_{-j,t}$ is $\gamma_{wj}$ sub-Gaussian. Also we assume $E\tilde{U}_{-j,t}\tilde{U}_{-j,t}^{\prime}=I_{p-1}$. Without losing any generality in the proofs we can use $\gamma_{w}$ constant instead of changing constants for each $j$, and $\gamma_{wj}$.
(ii). Define $f_{t}:=\Sigma_{f}^{1/2}\tilde{f}_{t}$, where $\tilde{f}_{t}:K\times 1$, is a subgaussian vector with $E[\tilde{f}_{t}\tilde{f}_{t}^{\prime}]=I_{K}$, and $\max_{1\leq k\leq K}|Ef_{tk}|\leq C<\infty$, where $C>0$ is a positive constant.
(iii). We assume $\|A\|_{\infty}\leq C<\infty$.
Assumption 4 is standard assumption used in Bing et al. (2021) Bunea et al. (2022). This assumption is needed for quadratic form high dimensional probability inequalities. It should be noted that we also relax the non-zero factor mean assumption used in earlier literature in Bing et al. (2021), and Bunea et al. (2022).
Define the uniform version of the signal to noise ratio
$\bar{\xi}=\min_{1\leq j\leq p}\xi_{j},$ (6)
where as in Bunea et al. (2022), they have a pointwise version
$\xi_{j}:=\frac{\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})}{\|\Sigma_{U,-j}\|_{2}},$ (7)
with $\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})$ representing the $K$ th largest eigenvalue of $A_{-j}\Sigma_{f}A_{-j}^{\prime}:p-1\times p-1$ matrix, and that is the smallest positive eigenvalue of that matrix since $p-1>K$, and $\Sigma_{f}:K\times K$ matrix.
We have a discussion about the role of the signal-to-noise ratio in a specific scenario. Since $A_{-j}:p-1\times K$ for all matrices $j=1,\cdots,p$, via Assumption 3, if the largest eigenvalue of $A_{-j}^{\prime}A_{-j}$ grows at the same rate $K$-th largest eigenvalue, we can assume $\lambda_{K}(A_{-j}^{\prime}A_{-j})\asymp d_{1n}^{2}$ and since we impose $max_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}=r_{n}$ by Assumption 3 then
$\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})\geq Eigmin(\Sigma_{f})\lambda_{K}(A_{-j}A_{-j}^{\prime})\geq c\lambda_{K}(A_{-j}^{\prime}A_{-j})\asymp d_{1n}^{2}.$
Utilizing (7) and relying on Assumption 3 we obtain a useful signal-to-noise ratio lower bound formulation
$\bar{\xi}\geq\frac{cd_{1n}^{2}}{r_{n}}.$ (8)
Depending on how fast/slow $d_{1n}$, and $r_{n}$ grow the signal-to-noise ratio lower bound in (8) can fluctuate between fairly small and fairly large quantities.
Assumption 5 below covers diagonal estimation of precision matrix. Note that $\bar{r},\bar{\Psi},\bar{\eta}$ are three random sequences that are explained in detail before Theorem 1, and they represent the complexity, prediction error, and prediction performance of the used technique respectively. In special cases of RRE, and adaptive PCR, we will show non-random upper bounds for these three key quantities.
To state Assumption 5 below, we find it convenient to introduce two rates, $r_{w1}$, $r_{w2}$ where $r_{w2}$ is the square root of the rate for off-diagonal terms in the precision matrix estimation. The rate $r_{w1}$ (compared to $r_{w2}$) is also used in estimation for the off-diagonal terms.
$r_{w1}:=\max\left(\sqrt{\frac{logn+\bar{r}}{n\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}}\frac{\bar{\Psi}}{\bar{\eta}}\frac{1}{\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}}\frac{1}{\bar{\eta}}}\right).$ (9)
$r_{w2}:=\max\left(r_{w1},\sqrt{\frac{K}{\bar{\xi}}}\right).$ (10)
Note the only difference between $r_{w1},r_{w2}$ is the second term in parentheses in (9)(10), since $\delta_{n}$ is nondecreasing in $n$, we can tell that $r_{w2}$ is the larger rate.
Define the relation between the factor loadings of $j$ th asset with all the others as $d_{2n}:=\max_{1\leq j\leq p}\|a_{j}^{\prime}\Sigma_{f}A_{-j}^{\prime}\|_{2}$.
Assumption 5.
(i). $\frac{d_{1n}}{\sqrt{p}}\frac{1}{\sqrt{K}}\max(1,\frac{\sqrt{logn}}{\sqrt{logp}})\to 0$.
(ii). $\sqrt{\frac{K^{3}}{n}}\to 0,\quad K\frac{\sqrt{max(logp,logn)}}{n^{1/2}}\to 0.$
(iii). $K\sqrt{\frac{p(logp)}{n}}r_{w1}\stackrel{{\scriptstyle p}}{{\to}}0$, and $d_{2n}r_{w2}\stackrel{{\scriptstyle p}}{{\to}}0$, and $d_{2n}$ is nondecreasing in $n$.
While the meaning of Assumption 5 will become clearer when we get to our final results, a few introductory remarks regarding the origin of the assumption are in place. First, Assumption 5(i) is a technical assumption that facilitates the proof of Theorem 2 for the finite sample results. [^4] One should note that this is a strengthened version of Assumption 3 when p and K are small; with larger p compared to n, Assumption 3 will suffice and this part of Assumption 5 will not be needed. Assumption 5(ii) relates to the tradeoff between number of factors and time span. In particular, it points out that the number of factors should be much smaller than the time span. Assumption 5(iii) involves stochastic terms and relates to consistency and rate of convergence of the off-diagonal terms estimates. As we will see later on, the role of Assumption 5(iii) will be discussed separately for adaptive PCR and RRE estimators.
3 Linear General Class of Estimators
We propose estimating the rows of the precision matrix of asset returns with our estimator. To show consistency of our estimators in high dimensional context we need to start with a precision matrix formula that can be tied to hidden factors. This is not a trivial task. One of the main obstacles is establishing relation between the matrix inversion and linear regression within hidden factor context. The difficulty is that formula for main diagonal terms and off-diagonal terms of the precision matrix is different and this type of approach tying hidden factors and general linear class of estimators with precision matrix estimation has not been done before. This is one of the contributions of our paper. Define the following parameter vector which will form a key term in precision matrix formula:
$\alpha_{j}^{*}:=argmin_{\alpha_{j}\in R^{p-1}}E[(y_{jt}-\mu_{j})-(Y_{-jt}-\mu_{-j})^{\prime}\alpha_{j}]^{2}.$
This definition shows that we can relate to linear regression context. Note that, the definition of $\alpha_{j}^{*}$ is not coming from the hidden factor model directly in (2)(3). We also need the following definitions before the next Lemma. Let $\Sigma_{f}=\Sigma_{f}^{1/2}\Sigma_{f}^{1/2}$ since $\Sigma_{f}$ is a positive definite matrix. Set $\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$, and $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j}$, and $\bar{G}_{j}:=I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}.$ Next, we show how they relate to hidden factor structure. Let $\Theta_{j,j}$ represent the main $j$ th diagonal term, and $\Theta_{j,-j}$ represent the $j$ th row without $j$ th element.
Lemma 1.
Under Assumptions 1-3, for $j=1,\cdots,p$
(i).
$\Theta_{j,j}=\frac{1}{\tau_{j}^{2}},\quad{\mbox{w}ith}\quad\tau_{j}^{2}=(\bar{a}_{j}^{\prime}\bar{a}_{j}+\sigma_{j}^{2})-\bar{a}_{j}^{\prime}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}.$
(ii).
$\Theta_{j,-j}=\frac{-\alpha_{j}^{*^{\prime}}}{\tau_{j}^{2}},\quad\mbox{with}\quad\alpha_{j}^{*}=\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}.$
(iii). Combining (i)(ii), we form $j$ th row of the precision matrix formula in terms of a hidden factor structure. Stacking each row will produce $\Theta$, the precision matrix.
Remark 1.
We make the following observations about the precision matrix formula and its relation to hidden factors:
  • One key result is that we will not impose arbitrary zero-exact sparsity restrictions on $\alpha_{j}^{*}$ as done in nodewise regression in Meinshausen and Bühlmann (2006), Callot et al. (2021).
  • Note that our precision matrix formula is not a direct extension nor an alternative representation of Sherman-Morrison-Woodbury matrix inversion formula. Rather, we start from a matrix algebra based formula and then impose hidden factor structure for outcomes to derive the precision matrix. The hidden factor structure is needed when we derive consistency of our estimator tied to the precision matrix formula.
  • Note that the formulas for $\alpha_{j}^{*}$ and $\tau_{j}^{2}$ are expressed in terms of the hidden factor based coefficients. This represents a novel contribution compared to existing literature—such as Callot et al. (2021), Caner et al. (2023), and Caner and Kock (2018)—which analyze settings without factors, consider errors in an observed factor model, or focus on regressors in a standard linear regression framework, respectively.
Connecting further to the existing literature, one should note that Bing et al. (2021) propose linear regression coefficient estimation and analyze the risk of an asset in a hidden factor model. However, neither the precision matrix estimation nor the above formulae are discussed therein. The main precision matrix estimation difficulty is that there are p rows which are tied to hidden factors. This leaves as widely open questions related to how and why linear regression estimators typically employed in hidden factor models should generalize to the precision matrix estimation. Lemma 1 and Theorem 1 below show how we manage to find a way to base the estimation on hidden factors but without actually imposing any identification conditions on hidden factors.
In that respect for each $j=1,\cdots,p$, we define
$\tilde{\alpha}_{j}:=\hat{B}(Y_{-j}\hat{B})^{+}y_{j}=\hat{B}(\hat{B}^{\prime}Y_{-j}^{\prime}Y_{-j}\hat{B})^{+}(\hat{B}^{\prime}Y_{-j}^{\prime}y_{j}),$ (11)
where $(.)^{+}$ is the Moore-Penrose inverse of a matrix, and $\hat{B}:p-1\times q$ with $q\leq p-1$ matrix, and the second equality is by (134) of Bunea et al. (2022). Dimension $q$ will be tied to specific estimators, and immediately below, we give examples about that. We let $\hat{B}$ be a deterministic matrix or a matrix that depends on $Y_{-j}$. Our estimator is based on analysis of precision matrix in Lemma 1. With different values of $\hat{B}$ we obtain different estimators. To see that, $\hat{B}=I_{p-1}$ is the ridgeless regression estimator (RRE) which is analyzed also after this section separately, so $q=p-1$. In case of adaptive PCR, $\hat{B}=V_{1,k_{j}}$ with $V_{1,k_{j}}$ being the $p-1\times k_{j}$ matrix where the columns are the eigenvectors of $Y_{-j}^{\prime}Y_{-j}/n$ matrix corresponding to the largest $k_{j}$ eigenvalues, and $q=k_{j}$, $j=1,\cdots,p$. We do not subscript $\hat{B}$ with $j$ to simplify the notation. Let us define the reciprocal of the main diagonal terms in the precision matrix based on (B.11) [^5] Note that only in case of RRE estimator as the subcase when $p>n$ the estimator simplifies and we benefit from that expression.
$\tilde{\tau}_{j}^{2}:=\frac{y_{j}^{\prime}(y_{j}-Y_{-j}\tilde{\alpha}_{j})}{n}.$ (12)
Then the $j$ the main diagonal term estimator in the precision matrix is:
$\tilde{\Theta}_{j,j}=1/\tilde{\tau}_{j}^{2},$ (13)
and the $j$ th row of the precision matrix estimate without $j$ th element (off-diagonal terms in the row) is, via (B.13)
$\tilde{\Theta}_{j,-j}=-\tilde{\alpha}_{j}^{\prime}/\tilde{\tau}_{j}^{2}.$ (14)
The $j$ th row of the estimate, to give an example, at $j=1$ will be $\tilde{\Theta}_{1}^{\prime}:=(\frac{1}{\tilde{\tau}_{1}^{2}},\frac{-\tilde{\alpha_{1}}^{\prime}}{\tilde{\tau}_{1}^{2}}).$
4 Uniform Consistency of General Estimator
In this section we provide first upper bounds on $l_{2}$ norm errors for each regression coefficient estimate, that is in (11). Next, we consider (12), and obtain upper bounds for the estimation errors. Analysis of (12) is key and we provide a new proof technique with using hidden factors without any exact sparsity assumptions. We use concentration inequalities for spectral norms, and high dimensional probability concepts to obtain upper bounds. Then we consider the consistency of each row of precision matrix estimate.
4.1 Bounding Estimation Errors: Regression Coefficients
In this section, our aim is to first obtain an upper bound for $\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2},$ which is essential in getting consistency of estimates of rows of the precision matrix. We explain three random terms that will be essential understanding our estimation error bounds. Let $P_{\hat{B}}$ be the projection matrix onto the range $\hat{B}$, i.e. $P_{\hat{B}}=\hat{B}[\hat{B}^{\prime}\hat{B}]^{+}\hat{B}^{\prime}=\hat{B}\hat{B}^{+}$, and the last equality is by Moore-Penrose matrix inverse rules such as (134) of Bunea et al. (2022). Define the rank of $Y_{-j}P_{\hat{B}}$ as
$\hat{r}_{j}:=rank(Y_{-j}P_{\hat{B}}).$ (15)
This term reflects the complexity, and increases the variance of the estimation of rows of the precision matrix, as we show below in Theorem 1. Define
$\hat{\eta}_{j}:=\sigma_{\hat{r}_{j}}^{2}(Y_{-j}P_{\hat{B}})/n,$ (16)
which is the $\hat{r}_{j}$- th largest singular value (squared, scaled by n) of $Y_{-j}P_{\hat{B}}$ ($n\times p-1$ matrix). This last term is obtained when we project $Y_{-j}$ onto range of $\hat{B}$, and will reduce the bias of estimating the rows of the precision matrix in Theorem 1 below. Next, we define
$\hat{\Psi}_{j}:=\sigma_{1}^{2}(Y_{-j}M_{\hat{B}})/n,$ (17)
where $M_{\hat{B}}:=I_{p-1}-P_{\hat{B}}$, and $Y_{-j}M_{\hat{B}}$ is $n\times p-1$ matrix. This term is the largest singular value of $Y_{-j}M_{\hat{B}}$ matrix scaled by $n$, and it represents the bias stemming from using a hidden factor structure to estimate $Y_{-j}$. This last term will increase the bias of estimating the rows of the precision matrix.
Define $\max_{1\leq j\leq p}\hat{r}_{j}:=\bar{r}$, $\min_{1\leq j\leq p}\hat{\eta}_{j}:=\bar{\eta}$, and $\max_{1\leq j\leq p}\hat{\Psi}_{j}:=\bar{\Psi}$. We have the following uniform bounds for estimators for off-diagonal terms in each row of the precision matrix
Theorem 1.
(i).Under Assumptions 1-4, $Klogn\leq cn$, with $c>0$ a positive constant, with probability at least $1-2/n^{C_{p}/2}$
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left(\frac{2logn+\bar{r}}{n\bar{\eta}}\right)+\frac{CK}{\bar{\xi}}\left(1+\frac{1}{\delta_{n}}+\frac{\bar{\Psi}}{\bar{\eta}\delta_{n}}+\frac{r_{n}}{\bar{\eta}\delta_{n}}\right).$
(ii) Under Assumptions 1-4, and $r_{n}/\delta_{n}\to 1$,
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}\left(\max(\frac{K}{\bar{\xi}},\frac{K\bar{\Psi}}{\bar{\xi}\bar{\eta}\delta_{n}},\frac{K}{\bar{\xi}\bar{\eta}},\frac{logn+\bar{r}}{n\bar{\eta}})\right)=O_{p}(r_{w2}^{2}),$
where $r_{w2}$ is defined in (9)-(10), by using $\delta_{n}$ is nondecreasing in $n$.
Remark 2.
  • One key issue is that our Theorem 1 forms a key building block for the consistency of rows of precision matrix estimation. The proof of this theorem uses a uniform concentration inequality type result that we show through a novel approach.
  • The first term on the right side of Theorem 1(i) is the variance term and the second term is the bias term.
  • There is no sparsity parameter on the upper bounds and rates.
  • Note that $\bar{\xi}$ plays an essential role in getting consistency. Larger signal-to-noise ratio is needed, especially if the number of factors $K$ grow fast.
  • The upper bounds given in the theorem are general. They do simplify as one moves to particular analyses of RRE and adaptive PCR subcases. However, it is useful to point out some generic principles. For example, $\bar{r}$ represents complexity, and affects the $l_{2}$ norm error adversely through the increase of the estimator’s variance. On the other hand, $\bar{\eta}$ is a positive influence, since a better prediction of asset returns gives a lower $l_{2}$ error. It should particularly be noted that this positive influence is reflected through the reduction of both the estimator’s bias and variance. $\bar{\Psi}$ shows how badly hidden factor estimates predict asset returns and it has an adverse effect on $l_{2}$ error. This term increases the bias of the estimator.
4.1.1 Adaptive PCR
In this subsection we consider principal components regression (adaptive PCR from now on) estimators. Let us designate $rank(Y_{-j})\geq k_{j}\geq 1$, where $k_{j}$ is a positive integer and changes with $j=1,\cdots,p$. The adaptive PCR estimators use $\hat{B}=V_{1,k_{j}}$, where $V_{1,k_{j}}$ is $p-1\times k_{j}$ matrix where $V_{1,k_{j}}$ represents a matrix which $k_{j}$ columns (eigenvectors) correspond to the largest $k_{j}$ eigenvalues of $Y_{-j}^{\prime}Y_{-j}/n$ for $j=1,\cdots,p$. The adaptive PCR estimator is, for $j=1,\cdots,p$
$\hat{\alpha}_{PCR,j}=V_{1,k_{j}}(Y_{-j}V_{1,k_{j}})^{+}y_{j},$
where $(.)^{+}$ is Moore-Penrose inverse of a matrix.
Before simplifying the terms in the upper bound in Theorem 1 we define $\hat{\lambda}_{k_{j}}:=\frac{1}{n}\sigma_{k_{j}}^{2}(Y_{-j})$, for example with $k_{j}=1$ we have the largest singular value of $Y_{-j}$, which is $\sigma_{1}^{2}(Y_{-j})/n$ (squared-scaled by $n$). One big issue is the choice of $k_{j}$. To choose $k_{j}$ we adapt the so-called elbow method, (for details, see (19) of Bing et al. (2021))
$\hat{s}_{j}:=\max\{k_{j}\geq 0:\hat{\lambda}_{k_{j}}\geq C_{0}\Delta_{-j}\},$ (18)
with $C_{0}>1$ and, $c_{\Delta}>1$ are positive constants above one,
$\Delta_{-j}:=c_{\Delta}[\|\Sigma_{U,-j}\|_{2}+tr(\Sigma_{U,-j})/n].$ (19)
This type of choice is done to minimize the risk of adaptive PCR in Bing et al. (2021). Here we need results uniformly over $j=1,\cdots,p$ and the choice given in (19) and above is a feasible estimator of $s_{j}$ after following Remark 3- fifth item to Corollary 1. Set $k_{j}=\hat{s}_{j}$. By definition of $\hat{s}_{j}$ above
$\hat{\lambda}_{\hat{s}_{j+1}}<C_{0}\Delta_{-j}\leq\hat{\lambda}_{\hat{s}_{j}}.$ (20)
We show in the proof of Corollary 1 in Online Appendix
$\max_{1\leq j\leq p}P(\hat{s}_{j}\leq K)\to 1.$ (21)
which will simplify the upper bound for $l_{2}$ norm error for our estimators. We prove the following error bounds for adaptive PCR based estimator for the off-diagonal terms in a given row for the precision matrix.
Corollary 1.
Under Assumptions 1-4 with $10<c^{\prime}<$ 1 a positive constant, $Klogn\leq cn$, with $c$ a positive constant, and $r_{n}/\delta_{n}\to 1$
(i).
$\max_{1\leq j\leq p}\|\hat{\alpha}_{PCR,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left[\frac{2logn+K}{n\delta_{n}}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{3}{\delta_{n}}\right],$
with probability at least $1-2/n^{C_{p}/2}-exp((c^{\prime}-1)n)$.
(ii).
$\max_{1\leq j\leq p}\|\hat{\alpha}_{PCR,j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}\left(max(\frac{logn+K}{n\delta_{n}},\frac{K}{\bar{\xi}})\right).$
Remark 3.
  • This is a new result for precision matrix row estimation. It clearly shows, that the estimator’s variance (i.e. the first square bracketed term on (i)-(ii)) goes to zero if $(logn+K)/n\delta_{n}\to 0$. Given $Klogn<cn$ is assumption, this is easily satisfied. It should be noted that large number of factors can be detrimental unless noise $\delta_{n}$ is large.
  • Regarding the bias term (the second square bracket in (i)) and the corresponding rate in (ii) the signal to noise ratio $\bar{\xi}$ should grow larger than $K$ to have consistency of the estimator. Utilizing, for example (8), we may need $Kr_{n}/d_{1n}^{2}\to 0$.
  • We note that with $p>n$, consistent estimation of $\alpha_{j}^{*}$ is possible, and this is also true with $p\leq n$ as well.
  • If we fix $p/n=\gamma\in[0,\infty)$ and then we increase $n,p$ together, and we keep $\gamma$ fixed to check consistency. In $(logn+K)/(n\delta_{n})$ condition, $\delta_{n}$ can grow with $n,p$ both increasing and have a positive effect on consistency. The price here is $logn$ rate, but that is not large compared to $n$ rate in the denominator. In the second condition, $K/\bar{\xi}$ is via (8) tied to $Kr_{n}/d_{1n}^{2}$ rate as discussed in second item in Remark 3 here. Increasing $n,p$ together can increase $r_{n}$, but also $d_{1n}$. Clearly, one observes a tradeoff in the $\delta_{n},r_{n}$ (both representing the noise) role in getting consistency. In particular, the noise can reduce variance but can increase bias. Note that $d_{1n}$ shows the strength of factor loadings $A_{-j}$. With $Kr_{n}/d_{1n}^{2}\to 0$ we have consistency.
[…]
4.1.2 Ridgeless Regression Estimation
In this part, we consider RRE when $p-1<n$ first and then analyze $p>n$ case. In that respect $\hat{B}=I_{p-1}$, which simplifies the upper bound in Theorem 1. We have
$\hat{\alpha}_{RRE,j}=(Y_{-j})^{+}y_{j},$
for $j=1,\cdots,p$. Note that when $p-1<n$, RRE estimation is equivalent to OLS, and by using (132) of Bunea et al. (2022), $\hat{\alpha}_{RRE,j}=(Y_{-j})^{+}y_{j}=(Y_{-j}^{\prime}Y_{-j})^{-1}Y_{-j}^{\prime}y_{j}.$ We show that after Corollary 2-Remark 4, indeed in an important subcase, OLS will not perform well in finite samples.
Corollary 2.
(RRE: $p-1<n$)
(i). Under Assumptions 1-4, and $Klogn<cn$, with $c>0$ a positive constant, with probability at least $1-1/n^{C_{p}/2}$, and $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left[\frac{2logn+p}{Cn}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{1}{\delta_{n}}+\frac{1}{C}\right].$
(ii). Under the same Assumptions in (i)
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}(max(\frac{logn+p}{n},\frac{K}{\bar{\xi}})).$
Remark 4.
  • Note that when $p$ is fixed , $n$ is fixed and $p/n=\gamma\in(0,1)$, where $\gamma$ is a positive fraction between 0 and 1, RRE or equivalently OLS does not perform well in finite samples. This is clear from the first term on the right side in Corollary 2(i). Even when $p/n\to\gamma$, with $p\to\infty,n\to\infty$, there is no consistency.
  • In large samples, a comparison with adaptive PCR shows that since $p>K$ by Assumption 3 we expect adaptive PCR to have a better rate in estimation error. Signal to noise ratio affects both estimators in the same way by reducing the bias term. Note that in the variance term in adaptive PCR, Corollary 1(ii), we may have $\delta_{n}\to\infty$ in the denominator so adaptive PCR has extra advantage of having a better variance than RRE ($p-1<n$). Adaptive PCR will perform better when $p/n=\gamma\in(0,1)$ case unlike RRE here asymptotically. RRE will be consistent when $p/n\to 0$.
  • Note that there is one less bias term in upper bound in RRE compared with adaptive PCR due to $\bar{\Psi}=0$ in RRE. This may be the only advantage in finite sample context.
Now we provide the Corollary for the RRE case when $p-1>n$. First we find it useful to introduce the effective rank as $re(\Sigma_{U,-j}):=\frac{tr(\Sigma_{U,-j})}{\|\Sigma_{U,-j}\|_{2}}.$ Next, we relax Assumption 3(i) that relates $p$ to $n$. We show that in RRE with $p-1>n$ case, we can allow $p$ to grow exponentially in $n$. This implies that we can have much larger, in dimension, precision matrix estimation in RRE compared with adaptive PCR.
Assumption $3^{*}(i)$. We allow $p\leq\frac{exp(C_{p}n/2)}{4}.$ Now we provide our coefficient estimation result that will be helpful for precision matrix estimation based on RRE.
Corollary 3.
(RRE: $p-1>n$)
Under Assumptions 1-2, and Assumptions $3^{*}(i)$-3(ii)-4 and $C>0$, and for all $j=1,\cdots,p$, $re(\Sigma_{U,-j})>Cn,$ for $C$ a positive constant $Klogn<cn$, with $c$ a positive constant, and $\tilde{U}_{-j}$ has independent entries, with $r_{n}/\delta_{n}\to 1$
(i).
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left[\frac{2logn+n}{Cn\delta_{n}}\right]+\frac{CK}{\bar{\xi}}\left[1+\frac{1}{\delta_{n}}+\frac{1}{C\delta_{n}}\right].$
with probability at least $1-exp(-C_{p}n/2)$, and
(ii).
$\max_{1\leq j\leq p}\|\hat{\alpha}_{RRE,j}-\alpha_{j}^{*}\|_{2}^{2}=O_{p}(max(\frac{logn+n}{n\delta_{n}},\frac{K}{\bar{\xi}})).$
Remark 5.
[…]
  • Note that we may see double descent in RRE for off-diagonal term estimation of the parameter vector $\alpha_{j}^{*}$ in the precision matrix. We fix $n$ and $K$ and analyze the changes in $p$. To see that in the previous Corollary we need $p/n\to 0$ to get estimation errors converging to zero. So very small $p$ in OLS (RRE $p-1<n$) case provides estimation error converging to zero. Then in minimum-norm interpolator case here (RRE with $p-1>n$), with $p/n\to\infty$ with $n$ fixed, and noise $\delta_{n}$ may grow with larger $p$ as in Assumption 2 $\frac{logn+n}{n\delta_{n}}\to 0,$ and $K/\bar{\xi}\to 0$ as discussed above since $r_{n},d_{1n}^{2}$ depend on $p$ and if $d_{1n}^{2}$ dominates $r_{n}$ we may see estimation errors converging to zero, and hence double descent can be seen. To see double descent Assumption $^{*}$(i),(ii) and (8) is helpful.
  • We compare now RRE in this case with adaptive PCR. In RRE there is one less bias term compared with adaptive PCR ($\bar{\Psi}=0$ in RRE) in finite sample upper bound. Since $K<n$, adaptive PCR has a smaller variance term compared with RRE. Asymptotically bias terms have the same rate $K/\bar{\xi}$. So asymptotically adaptive PCR have smaller estimation errors, but in finite samples due to one less bias term in RRE, RRE may have smaller bias, compared to adaptive PCR.
  • Adaptive PCR also may have one more theoretical advantage compared with RRE. We need the noise $\delta_{n}$ to be increasing in $n$ to get consistency in RRE, whereas in adaptive PCR this is not needed. But in financial markets we expect $\delta_{n}\to\infty$ so this may not be of an issue between two estimators in financial markets.
4.2 Estimation of Diagonal Precision Matrix Entries
We have the following Theorem for the reciprocal of the main $j$-th diagonal term in the precision matrix. The technique in Bing et al. (2021) considers only simple linear regression, while the precision matrix estimation and especially a diagonal term estimation are not handled. The proof of Theorem 2 extends the literature on diagonal term estimation of a large precision matrix. The results and the proof are novel and show the rate of convergence and the non-asymptotic upper bound for estimation errors, as well as the consistency of the estimator. Note that $r_{w1},r_{w2}$ rates are defined just before Assumption 5 in the main text.
Theorem 2.
(i). Under Assumptions 1-4 5(i), with $r_{n}/\delta_{n}\to 1$,
$\max_{1\leq j\leq p}|\tilde{\tau}_{j}^{2}-\tau_{j}^{2}|$ $=$ $O_{p}\left(K(\sqrt{\frac{p(logp)}{n}})r_{w1}\right)$
$+$ $O_{p}\left(d_{2n}r_{w2}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\max(\sqrt{logp},\sqrt{logn})}{n^{1/2}})\right).$
(ii). Under Assumptions 1-5, with $r_{n}/\delta_{n}\to 1$,
$\max_{1\leq j\leq p}|\tilde{\tau}_{j}^{2}-\tau_{j}^{2}|=o_{p}(1).$
(iii). Under Assumptions 1-5, with $r_{n}/\delta_{n}\to 1$, where $c>0$ is a positive constant
$\min_{1\leq j\leq p}\tau_{j}^{2}\geq c-o(1).$
Remark 6.
  • We show that estimation of the reciprocal of the main diagonal terms in precision matrix estimation depends on number of factors, $K$, sample size-time span $n$, number of assets $p$, and signal-to noise ratio $\bar{\xi}$ as well as three terms $\bar{r},\bar{\Psi},\bar{\eta}$ which represents complexity, prediction errors, and prediction performance respectively. These last three terms are random and will be upper bounded with non-random terms, wpa1, in subcases of RRE and adaptive PCR below in Corollaries A.1-A.3.
  • When $p>n$ and $p/n\to(1,\infty)$ it is possible to get consistency since $p$ only enters the estimation errors directly via $\sqrt{logp}$ and with a larger n, the estimation errors will go to zero, if $K\sqrt{logp}\,r_{w1}=o_{p}(1)$, and $K\sqrt{logp/n}=o_{p}(1)$.
  • Number of factors $K$ play a key role, to have Assumption 5(ii), we need $K=o(n^{1/3})$, and there are more restrictions on $K$ in Assumptions 5(i), (iii). Larger number of factors provide a real roadblock in getting consistency.
  • Most important is the fact that estimation of the reciprocal of the main diagonal terms in the precision matrix. This is more difficult compared with estimating each row of the precision matrix (without the main diagonal term). Theorem 1(ii) clearly shows that the rate of estimation error for off-diagonal-row elements is $r_{w2}$, but our Theorem 2 shows for the reciprocal of the main diagonal term, this is more complicated and slower than $r_{w2}$ since via Assumption 5(iii), we need $d_{2n}r_{w2}=o_{p}(1)$ for consistency in addition to other part of Assumption 5(iii) as well as Assumption 5(ii).
  • Note that Theorem 2 excludes the case of RRE with $p-1>n$ case, since formula in (12) is equal to zero due to interpolation. We provide an alternative estimator for RRE with $p-1>n$ case. That case is proved separately in Corollary OA.3. Subcases of adaptive PCR and RRE with $p-1<n$ cases are shown in Corollaries OA.1-OA.2.
4.3 Consistent Estimation of Rows
We provide our theorem for estimating rows of the precision matrix. This shows that each row of the precision matrix in general case can be estimated, uniformly over all rows, consistently through our estimator. Specific results regarding adaptive PCR and RRE will be discussed in subsections 4.3.1-4.3.2. Note that the main diagonal terms rate determine the rate of convergence. The off-diagonal terms estimation is consistent but converges at a faster rate than the main diagonal terms. This can be seen through comparing the rates in Theorem 1 and via proof of Theorem 3 and Remark-Item 4 in Theorem 2. Denote each row of $\Theta$ as $\Theta_{j}^{\prime}$, which is a $1\times p$ vector. As an example, first row can be represented as $\Theta_{1}^{\prime}:=(\frac{1}{\tau_{1}^{2}},\frac{-\alpha_{1}^{*^{\prime}}}{\tau_{1}^{2}})$. Let $\Theta_{j}$ be the column representation of the row $\Theta_{j}^{\prime}$. Let the row estimator be $\tilde{\Theta}_{j}^{\prime}$, and for the first row this is $\tilde{\Theta}_{1}^{{}^{\prime}}:=(\frac{1}{\tilde{\tau}^{2}},\frac{-\tilde{\alpha}_{1}^{{}^{\prime}}}{\tilde{\tau}_{1}^{2}})$.
Theorem 3.
Under Assumptions 1-5, with $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}\|\tilde{\Theta}_{j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $=$ $O_{p}(K\frac{\sqrt{plogp}}{\sqrt{n}}r_{w1})+O_{p}(d_{2n}r_{w2})$
$+$ $O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\frac{\sqrt{\max(logp,logn)}}{n^{1/2}})\right)=o_{p}(1),$
where $r_{w1},r_{w2}$ are defined in (9)(10).
Remark 7.
  • First, we see that if $p/n\to[0,\infty)$, we obtain consistency and the price that we pay in terms of number of assets, $p$ is just $\sqrt{logp}$ at the first glance, as can be seen in Assumption 5. However, we have terms such as $\bar{r},\bar{\Psi},\bar{\eta}$ which may depend on $p$, and there will be analysis of these terms in adaptive PCR, and RRE cases.
  • We see that through definition of $r_{w1},r_{w2}$ one of the keys to consistent estimation is the ratio $K/\bar{\xi}$, and this ratio needs to go to zero. So we cannot have a large hidden factor system and expect consistency. Key is the tradeoff between two terms. With low signal-to-noise ratio, the number of hidden factors has to be small.
  • Note that the rate for Theorem 3 is obtained for all linear estimators in Section 3, except RRE when $p-1>n$, the interpolating estimator. This is explained in detail, below in Corollary 6 and Appendix Corollary OA.3.
4.3.1 Adaptive PCR Estimator
In case of adaptive PCR, the random terms in the upper bound in Theorem 3, can be bounded as: $\bar{r}\leq K$ wpa1, $\bar{\Psi}/\bar{\eta}\leq 1$ wpa1, $\bar{\eta}\geq\delta_{n}$ with wpa1. So Theorem 3 rate can be simplified in adaptive PCR, with $\hat{\Theta}_{PCR,j}^{\prime}$ representing the estimator of $j$ th row of $\Theta$. The following Corollary follows from Theorem 3 by replacing $r_{w1},r_{w2}$ with $r_{w1,PCR},r_{w2,PCR}$ rates that are shown before Corollary OA.1 in Online Appendix.
Corollary 4.
Let $p/n\to[0,\infty)$ and with Assumptions 1-5, and with $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}\|\hat{\Theta}_{PCR,j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $=$ $O_{p}\left(K\sqrt{logp}\max(\frac{\sqrt{logn+K}}{\sqrt{n\delta_{n}}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}\delta_{n}}})\right)$
$+$ $O_{p}\left(d_{2n}\max(\frac{\sqrt{logn+K}}{\sqrt{n\delta_{n}}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})\right)$
$+$ $O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\frac{\sqrt{\max(logp,logn)}}{n^{1/2}})\right)=o_{p}(1).$
Remark 8.
  • The price of consistency is $\sqrt{logp}$ by number of assets, even when $p>n$. So with $p/n\to(1,\infty)$, to get consistency we need the signal-to-noise ratio dominate the number of factors.
  • We analyze how consistency can be achieved in this adaptive PCR. We start with Assumption 5(iii) which may be the most stringent term for consistency. We use example (8) in Assumption 5(iii) explanation to illustrate our point. Take the case of $p/n\to\gamma$, with $\gamma\in(1,\infty)$, $\gamma$ is a positive constant. Case of $p/n\to\gamma$, with $\gamma\in[0,1]$ is similar hence skipped. What if we increase $n,p$ together but keep $\gamma$ fixed. Consistency will require,
    $\frac{K\sqrt{log(\gamma n)}\sqrt{logn+K}}{\sqrt{n\delta_{n}}}\to 0,$ (22)
    which is possible with Assumption 5(ii) imposing $K=o(n^{1/3})$, and since $\delta_{n}$ is nondecreasing in n . Then we need
    $\frac{K^{3/2}\sqrt{log(\gamma n)}}{\sqrt{\delta_{n}\bar{\xi}}}\to 0$ (23)
    which can be possible with (8), $r_{n}/\delta_{n}\to 1$ and since $\frac{K^{3/2}\sqrt{log(\gamma n)}}{d_{1n}}\to 0,$ with increasing $d_{1n}$ with $n$, and can grow larger than $K^{3/2}\sqrt{logn}$. Next, we consider in Assumption 5(iii) $d_{2n}\sqrt{\frac{logn+K}{n\delta_{n}}}\to 0,$ since $K=o(n^{1/3})$ and $d_{2n}$ can diverge with n but not too fast (i.e. such as a slowly varying. function in n). Then for the second part of that consistency condition we also need with example (8) $d_{2n}\sqrt{\frac{K}{\bar{\xi}\delta_{n}}}\leq\frac{d_{2n}K^{1/2}}{d_{1n}}\to 0.$ So the signal $d_{1n}$ can be large and bring consistency.
  • An interesting connection of the above theorem’s results and Fan et al. (2013) can be made as well. First, one recalls that Fan et al. (2013) estimates hidden factors through principal components and feeds such estimates into covariance matrix of returns which is then inverted via Sherman-Woodbury- Morrison formula. However, underlying hidden factor rotation results in having favorable rates of convergence provided that p is large.
4.3.2 Ridgeless Estimators
When we consider the sub case of RRE with $p<n-1$, we set wpa1 $\bar{r}<p$, $\bar{\Psi}=0$, and $\bar{\eta}\geq c$, where $c$ is a positive constant in Theorem 3 result above. Let the $j$ th row of estimator of the precision matrix be $\hat{\Theta}_{RRE,j}^{\prime}$. To obtain Corollary 5, we use Corollary OA.2 and Corollary 2 with $r_{w1},r_{w2}$ in Theorem 3 replaced with $r_{RRE-l}$ rate after Corollary OA.1 in Online Appendix.
Corollary 5.
Set $p/n=a_{2n}\to 0$, and under Assumptions 1-5, and $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}\|\hat{\Theta}_{RRE,j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $=$ $O_{p}\left(\max(K\sqrt{a_{2n}logp},1)\max(\frac{\sqrt{logn+p}}{\sqrt{n}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})\right)$
$+$ $O_{p}\left(d_{2n}\max(\frac{\sqrt{logn+p}}{\sqrt{n}},\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})\right)$
$+$ $O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\frac{\sqrt{logn}}{n^{1/2}})\right)=o_{p}(1).$
Remark 9.
  • Note that unless $p/n\to 0$, it is not clear that RRE will be consistent. This is a major handicap compared with adaptive PCR, since adaptive PCR is consistent even when $p/n=\gamma\in(0,1)$. Also we see that with $K<p$, the rate of convergence of adaptive PCR in case of $p/n\to 0$ will be better than RRE here. Only in finite samples RRE, may have an advantage due to nuisance parameter $\bar{\Psi}=0$ in RRE. RRE can work well only when $p$ is very small compared with $n$. Also, when $p-1<n$ there is a notable advantage of RRE. Namely, it can be run without tuning parameter as RRE is equivalent to OLS estimator in this case.
  • One remaining question is with $p/n=a_{2n}\to 0$ then which method will have a better rate. Consider the first part of Assumption 5(iii), then in RRE
    $a_{2n}^{1/2}K(\sqrt{logp})r_{RRE-l}=a_{2n}^{1/2}K\sqrt{logp}\max(\sqrt{\frac{logn+p}{n}},\sqrt{\frac{K}{\bar{\xi}}}),$
    and in adaptive PCR the same rate is
    $a_{2n}^{1/2}K(\sqrt{logp})r_{w1,PCR}=a_{2n}^{1/2}K\sqrt{logp}\max(\sqrt{\frac{logn+K}{n\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}}),$
    which will be smaller since $K<p$ and $\delta_{n}$ is nondecreasing in n. The second part of Assumption 5(iii) follows a similar pattern and hence adaptive PCR is a better estimator asymptotically than RRE when $p-1<n$.
Next, we consider RRE estimator when $p-1>n$. Note that in this scenario main diagonal term estimation is done differently since RRE is an interpolating estimator. We prove this after Corollary OA.2. In the proof of Theorem 3, it is clear that the rate of convergence should be the slower of the main diagonal or off-diagonal terms. Let the j-th row of the precision matrix be $\hat{\Theta}_{INT,j}$. Corollary 6 below proves $\hat{\Theta}_{INT,j}$’s consistency by combining Corollaries 3 and OA.3.
Corollary 6.
Under Assumptions 1-2, Assumption $3^{*}(i)$ Assumptions 3(ii)-4-5(ii), with signal-to-noise condition $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$, and $\min_{1\leq j\leq p}re(\Sigma_{U_{-j}})>Cn$ and $\tilde{U}_{-j}$ matrix with independent entries,
$\max_{1\leq j\leq p}\|\hat{\Theta}_{INT,j}^{\prime}-\Theta_{INT,j}^{\prime}\|_{2}$ $=$ $O_{p}\left(\max(\frac{\sqrt{logn+n}}{\sqrt{n\delta_{n}}},\frac{\sqrt{K}d_{2n}}{\sqrt{\bar{\xi}}})\right)$
$+$ $O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\frac{\sqrt{logp)}}{n^{1/2}})\right)=o_{p}(1).$
Remark 10.
  • First of all we still get consistency with $p-1>n$, as long as signal-to-noise ratio dominates number of factors multiplied by $d_{2n}$. This can be seen easily by $K\sqrt{logp/n}\to 0$, and also we need signal-to-noise condition: $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$.
  • It is not clear from the rates of convergence that RRE or adaptive PCR may do better. But it is possible that with low number of factors, adaptive PCR can perform better. But there is no tuning parameter estimation in RRE.
  • As described in Remark 5-item 2 of Corollary 3 and item 1 above, double descent is possible in RRE in precision matrix estimation. This extends the concept of double descent from linear regression to precision matrix estimation. This may have important consequences in empirics. We show one of these findings in Sharpe Ratio in US Stock market.
4.4 Uses of Precision Matrix Estimate
Here in this section we discuss the usage of our precision matrix estimator in three different cases of interest in economics and finance. In addition to the examples that we will discuss below, we should also add that many of our results are of independent statistical interest as well. We single out two particular features of our statistical analyses, (i) lack of exact sparsity assumption between outcomes with model-free dense estimators and (ii) an explicit connection of the estimation error to signal-to-noise ratio, as results that significantly advance the current state of the art in the relevant statistics precision matrix estimation related literature.
In financial econometrics and finance, one of the key issues is large portfolio formation. Define the number of assets of a large portfolio as $p$, and the sample size as $n$. Two precision matrix estimating approaches distinguished themselves in recent literature: (i) the first one that relies on the estimation and inversion of the covariance matrix (see, e.g., Fan et al. (2013) and Ledoit and Wolf (2017) where covariance matrix is obtained via hidden factor model and shrinkage estimation, respectively); and (ii) the second one that directly estimates the precision matrix and relies on the so-called nodewise regression approach (see, e.g., Callot et al. (2021) and Caner et al. (2023) for nodewise regression with observed factors). We introduce a new estimator that can be used to consistently estimate the Sharpe Ratio even when $p>n$ (Sharpe Ratio is one of the key practical financial metrics as it can measure return-risk ratio for a given portfolio).
The second possibility is a simple modification of our method can be used to estimate inverted Gram matrix thereby replacing the so-called relaxed inverse typically used to get a high dimensional linear regression parameters confidence intervals (see, e.g., van de Geer et al. (2014)). Our idea can also be used to form and change the debiased high dimensional estimators used in van de Geer et al. (2014) and Caner and Kock (2018). Below, we provide some details on how this can be achieved.
The third possibility for the utilization of our methodology is the network analysis. One first recalls that nodewise regression can be tied to network analysis (see, e.g., Meinshausen and Bühlmann (2006). Given the connection between our estimator and nodewise regression, one then observes that our estimator fits the same network structure but with an advantage of not needing to impose sparsity on the links between two nodes.
We have three examples of our estimators, and the first one is Sharpe-Ratio analysis, and the second one is the de-biased high dimensional estimator formation for the confidence intervals for structural parameters in cross section analysis, and the third one is network analysis. In this part, first, we provide a Sharpe Ratio metric that use precision matrix of asset returns. The metric is the Sharpe ratio (squared) from the maximum Sharpe Ratio portfolio. There are other metrics as discussed heavily in Caner et al. (2023) for example, but the emphasis of our paper is introduction of a new class of estimators for the precision matrix not the detailed analysis of financial metrics. We also think that the analysis of financial metrics can be done in another paper with this approach, since the proofs will be non-standard and will take a larger space with very detailed analysis through several matrix norm inequalities. The metric below is chosen since the contribution of the precision matrix is very clear to the financial metric. To obtain the (squared) Sharpe ratio from the maximum Sharpe Ratio portfolio we maximize
$w_{MS}^{*}:=argmax_{w\in R^{p}}w^{\prime}\mu\quad{\mbox{s}ubject}\quad{\mbox{t}o}\quad w^{\prime}\Sigma w\leq\sigma^{2},$
where $\sigma^{2}$ is the variance of the portfolio. The optimal solution is $w_{MS}^{*}=\frac{\sigma\Sigma^{-1}\mu}{\sqrt{\mu^{\prime}\Sigma^{-1}\mu}}.$ Substituting this into Sharpe Ratio formula we have $SR_{MS}=\frac{w_{MS}^{*^{\prime}}\mu}{w_{MS}^{*^{\prime}}\Sigma w_{MS}^{*}}=\sqrt{\mu^{\prime}\Sigma^{-1}\mu}.$ The estimator is: $\widehat{SR}_{MS}^{2}:=\hat{\mu}^{\prime}\tilde{\Theta}^{\prime}\hat{\mu},$ with $\hat{\mu}:=\frac{1}{n}\sum_{t=1}^{n}Y_{t}$.
This can be found in Caner et al. (2023) as well as several other papers in the literature, and used in practice. The prevalent approaches in present day literature impose sparsity either on the covariance or on the precision matrix of errors, and prove consistency results relying on those assumptions. Without exact sparsity, the proofs will not be valid for $p>n$ case. On the other hand, our methodology allows to circumvent the sparsity assumption. The second issue is that rates of convergence are usually only tied to the number of assets, p, sample size, n, and the number of factors, K. Our methodology allows to additionally tie them to signal-to-noise ratio as well.
In our second example, we show how our precision matrix can be modified-used to form a debiased lasso estimator for estimating structural parameters, and this type of estimator can then be used to have confidence intervals for these structural parameters. Our setup is a simple linear high dimensional linear regression as in Caner and Kock (2018) and van de Geer et al. (2014).
$Z=Y\beta+\epsilon,$
where $Z$ represents the outcome variables here, and it is $n\times 1$ vector, where $Z_{i},i=1,\cdots,n$ are cross-section units (instead of $t$ index in the other parts of the paper, we use $i$ for cross section data), and $Y:n\times p$ matrix, where each row represents cross section units, and $j=1,\cdots,p$ represent regressors. Let $\beta:p\times 1$ be the structural parameters of interest, and let $s_{0}$ of them to be nonzero. We allow $s_{0}$ n to be nondecreasing in $n$. Let $\epsilon:n\times 1$ represent the errors. Also let $EY_{i}\epsilon_{i}=0$, and $\epsilon_{i}$ is iid as well as $Y_{i}$. For example, $Z$ can represent average (over time) growth rate of GDP of a $n$ group of countries, and each column of $Y$ represents each regressor, such as public debt to GDP ratio, current account deficit, inflation, that contributes to this average GDP growth rate. We also assume existence of the following relation
$Y=FA+U,$
as can be formed by (4)(5), so that regressors are governed by latent factors. Note that using lasso estimation of $\beta$ results in biased estimates and is not uniformly consistent, hence any confidence intervals will not be valid and not asymptotically honest as discussed in Caner and Kock (2018). To solve this issue, a method is debiasing the lasso estimators. However, debiased lasso estimators depend on the estimation of the inverse of the population inner-product $EY_{i}Y_{i}^{\prime}$, where $Y_{i}:p\times 1$ vector, (row of $Y$ represented in column form). It is not possible to use sample $Y^{\prime}Y/n$ and invert it to get an estimate since if $p>n$, $Y^{\prime}Y/n$ is singular. First we show how to form $[EY_{i}Y_{i}^{\prime}]^{-1}$ which is a $p\times p$ matrix. This can be formed exactly with the analysis in Appendix B, (with only imposing $\mu_{j}=0,\mu_{-j}=0_{p-1}$, where $0_{p-1}$ vector is a p-1 vector of zeroes in each element). Also see p.157 of Caner and Kock (2018) for this as well. Our estimator for $\tilde{\Theta}$ is exactly the same as in Section 4 with no change. Then the debiased estimator, $\hat{b}$, can be formed through, using the analysis in Section 3.1, equations (7)-(9) of Caner and Kock (2018) but with our new estimator of inverse of $EY_{i}Y_{i}^{\prime}$ rather than the nodewise regression in Caner and Kock (2018),
$\hat{b}=\hat{\beta}+\tilde{\Theta}Y^{\prime}(Z-Y\hat{\beta})/n,$
where $\hat{\beta}$ is the lasso estimator for $\beta$ by regressing a $l_{1}$ penalized regression of $Z$ on $Y$. We want to focus on this estimator and how to form confidence intervals for our next project.
Our network estimator can be based on the theory in Appendix B, and it can be used, a detailed analysis is beyond the scope of this paper. But a start is the social interaction matrix estimation in De Paula et al. (2024).
5 Empirical Application
The objective of the empirical application is to conduct an out-of-sample comparison among estimations. We estimate Maximum Sharpe Ratio (MSR) portfolio using estimators, and use four commonly used financial metrics: the Sharpe Ratio (SR), average return, portfolio turnover, and portfolio variance. We use data from January 1995 to December 2023, with an out-of-sample period from January 2015 to December 2023. We selected all stocks included in the S&\&&P 500 for at least one month during the out-of-sample period, ensuring the data were available for the entire 1995-2023 period, resulting in 357 stocks. The out-of-sample period includes stock market of 2022 when S&\&&P 500 index was down by 18.1 %.
5.1 Metrics
We use a rolling window set up for the out-of-sample estimation of the Sharpe Ratio following Callot et al. (2021) and Caner et al. (2023). The sample size n is divided into in-sample $(1:n_{I})$ and out-of-sample $(n_{I}+1:n)$. Initially, we estimate the portfolio $\hat{w}_{n_{I}}$ in the in-sample period and the out-of-sample portfolio returns $\hat{w}^{\prime}_{n_{I}}y_{n_{I}+1}$. Then, we roll the window forward by one element $(2:n_{I}+1)$, forming a new in-sample portfolio $\hat{w}_{n_{I}+1}$ and caculating out-of-sample portfolio returns $\hat{w}_{n_{I+1}}^{\prime}y_{n_{I}+2}$. This process is repeated until the end of the sample.
The out-of-sample average return and variance without transaction costs are computed as follows: $\hat{\mu}_{os}=\frac{1}{n-n_{I}}\sum_{t=n_{I}}^{n-1}\hat{w}_{t}^{\prime}y_{t+1},\hat{\Sigma}^{2}_{y,os}=\frac{1}{n-n_{I}-1}\sum_{t=n_{I}}^{n-1}(\hat{w}_{t}^{\prime}y_{t+1}-\hat{\mu}_{os})$. We estimate the Sharpe Ratios without transaction costs: $SR=\frac{\hat{\mu}_{os}}{\hat{\Sigma}_{y,os}}$
When transaction costs are present, the definition of the metrics is modified. Let $c$ be the transaction cost, which we set 10 basis points as. Let $y_{P,t+1}=\hat{w}_{t}^{\prime}y_{t+1}$ be the return of the portfolio $P$ in period $t+1$; in the presence of transaction costs, the return is defined as $y^{Net}_{P,t+1}=y_{P,t+1}-c(1+y_{P,t+1})\sum_{j=1}^{p}|\hat{w}_{j,t+1}-\hat{w}^{+}_{j,t}|$ where $\hat{w}^{+}_{j,t}=\frac{\hat{w}_{j,t}(1+y_{j,t+1})}{1+y_{P,t+1}}$, with $y_{j,t+1}$ representing the excess return added to risk-free-rate for $j$th, and $y_{P,t+1}$ representing the portfolio $P$ excess return added to the risk-free rate. The adjustment made in $\hat{w}^{+}_{j,t}$ is because the portfolio at the end of the period has changed compared to the portfolio at the beginning of the period. Then, we define the mean and variance of the portfolio returns with transaction costs as: $\hat{\mu}_{os-c}=\frac{1}{n-n_{I}}\sum_{t=n_{I}}^{n-1}y^{Net}_{P,t}\mbox{ and }\hat{\Sigma}^{2}_{y,os-c}=\frac{1}{n-n_{I}-1}\sum_{t=n_{I}}^{n-1}(y^{Net}_{P,t}-\hat{\mu}_{os-c})^{2}$. The portfolio returns are replaced by the returns with transaction costs when we calculated the Sharpe Ratio with transaction costs. Finally, portfolio turnover is defined as: $\text{turnover}=\sum_{j=1}^{p}|\hat{w}_{j,t+1}-\hat{w}_{j,t}^{+}|.$
5.2 Methods
We conduct a comprehensive empirical analysis by fixing in-sample $n_{I}=240$ and considering both low-dimensional and high-dimensional cases across various portfolio sizes. Specifically, we selected 10, and 200 stocks from the pool to estimate the models with the same stocks for all windows in the low-dimensional experiments. For the high-dimensional experiment, we choose 357 (including all available stocks). We also analyzed portfolios with 25, 50, and 300 stocks, however, since these results are similar, they are not included here. We have 100 for low dimensional experiment and 325 stock high dimensional experiment in Online Appendix.
The model comparisons include several approaches from the literature. We consider the nonlinear shrinkage methods of Ledoit and Wolf (2017)(NL-LW) and the single-factor nonlinear shrinkage (SF-NL-LW), POET method from Fan et al. (2013), and the nodewise (NW) method from Callot et al. (2021), along with the feasible residual-based nodewise approach from Caner et al. (2023). SF-NL-LW results are very similar to NL-LW in terms of standard deviation, so we have not included that in our Table. But these are available from authors on demand. In particular, RNW-SF, RNW-3F, and RNW-5F refer to residual-based nodewise regressions using a single-factor model (market as the only factor), three-factor model, and five-factor Fama and French model, respectively. Our new method introduces several variations: RRE, PCR-1F, PCR-3F, PCR-5F, PCR-7F and PCR-Adaptive. These denote the use of a RRE estimator, Principal component with fixed 1, 3, 5, 7 hidden factors, and an adaptive method to select the number of hidden factors, respectively. We use these number of factors since they are used in the empirics in the literature and seven factor is used to show large number of factors.
5.3 Results
We report the results for Maximum Sharpe Ratio portfolio, with and without transaction costs. We have also done analysis for Global Minimum Variance portfolio, and the results are very similar. Results for the low-dimensional cases are presented in Tables 1 - 2, while high-dimensional case is shown in Table 3. In additionally, we use the same data but with different out-of-sample from January 2005 to December 2023 ($n_{I}=120$, $n-n_{I}=228$). This subsample yields similar results and is available upon request. In MSR portfolio, our PCR-based method is the best model across all 2 small-to-medium (low-dimensional) portfolios scenarios for Sharpe Ratio with transaction cost. Our RRE achieves the highest Sharpe Ratio for the largest size of portfolio with $p=357$. Let us give an example by examining three portfolios of varying sizes for MSR: a small low-dimensional portfolio, a moderate-sized portfolio, and a large-sized portfolio. In Table 1, for a portfolio of 10 stocks ($p=10,n_{I}=240$), PCR-1F is the best Sharpe Ratio with 0.2260. In Table 2 of 200 stocks ($p=200,n_{I}=240$), the highest Sharpe Ratio is with our PCR-7F at 0.2682 and the second best is our PCR-3F at 0.2263. Lastly, in Table 3, for the largest portfolio with all 357 available stocks, the highest Sharpe ratio is RRE at 0.2203.
We use Ledoit and Wolf (2008) test with circular bootstrap to test the significance of the Sharpe Ratio with transaction cost-winner. Specifically,
$H_{o}:SR_{winner}\leq SR_{0}\mbox{ vs }H_{a}:SR_{winner}>SR_{0},$ (24)
Where $SR_{winner}$ is the method that has the highest Sharpe Ratio in each table, testing against all other methods and significance is set at threshold of $p=0.05$. A """""*"" ∗ " indicates statistical significance from the winner.
To see whether a double ascent in the Sharpe Ratio is possible in RRE, we analyze a figure. Figure 1 shows the MSR Sharpe Ratio with transaction cost. In figure, we include the performance of RRE, Nodewise and PCR-3F, PCR-adaptive methods. We observe a double ascent pattern in the RRE estimator’s performance in terms of the Sharpe Ratio. For example, with the in-sample window size set at $n_{I}=240$, in low-dimensional case, the Sharpe Ratio starts at 0.2065. As the portfolio size increases to 200 stocks, the Sharpe Ratio declines to -0.0471. In the high-dimensional case, where the number of stocks exceeds 240, the SR increases to 0.2203 for a portfolio of 357 stocks. From a practitioner point of view, we observe that we can form a large portfolio with RRE, much larger than the sample size n, without any tuning parameters to fix as in other methods and get a good result.
Figure 1: MSR Sharpe Ratio vs. p
Table 1: Monthly Portfolio Performance of 10 Stocks, $n_{I}=240$, $n-n_{I}=108$
Estimator Maximum Sharpe Ratio
Return SD Sharpe Turnover
without transaction cost
RNW-GIC-SF 0.0101 0.0526 0.1913 0.0953
RNW-GIC-3F 0.0101 0.0519 0.1945 0.0889
RNW-GIC-5F 0.0102 0.0510 0.2002 0.0876
NW-GIC 0.0106 0.0505 0.2093 0.0888
POET 0.0104 0.0507 0.2051 0.0700
Ledoit-Wolf 0.0100 0.0522 0.1914 0.0927
Inter-RRE 0.0104 0.0530 0.1957 0.0929
PCR-1F 0.0109 0.0490 0.2221 0.0518
PCR-3F 0.0105 0.0492 0.2142 0.0576
PCR-5F 0.0106 0.0501 0.2125 0.0948
PCR-7F 0.0105 0.0507 0.2075 0.0795
PCR-Adaptive 0.0108 0.0489 0.2211 0.0514
with transaction cost
RNW-GIC-SF 0.0105 0.0526 0.1993 NA
RNW-GIC-3F 0.0106 0.0518 0.2040 NA
RNW-GIC-5F 0.0107 0.0509 0.2094 NA
NW-GIC 0.0110 0.0504 0.2184 NA
POET 0.0112 0.0502 0.2235 NA
Ledoit-Wolf 0.0105 0.0521 0.2025 NA
Inter-RRE 0.0109 0.0528 0.2065 NA
PCR-1F 0.0111 0.0491 0.2260 NA
PCR-3F 0.0108 0.0493 0.2188 NA
PCR-5F 0.0111 0.0500 0.2209 NA
PCR-7F 0.0109 0.0507 0.2160 NA
PCR-Adaptive 0.0110 0.0491 0.2249 NA
Table 2: Monthly Portfolio Performance of 200 Stocks, $n_{I}=240$, $n-n_{I}=108$
Estimator Maximum Sharpe Ratio
Return SD Sharpe Turnover
without transaction cost
RNW-GIC-SF 0.0053 0.0442 0.1205 0.2761
RNW-GIC-3F 0.0067 0.0447 0.1496 0.3171
RNW-GIC-5F 0.0072 0.0484 0.1483 0.3996
NW-GIC 0.0093 0.0429 0.2169 0.1517
POET 0.0057 0.0430 0.1336 0.1989
Ledoit-Wolf 0.0048 0.0521 0.0928 0.7320
Inter-RRE -0.0018 0.1168 -0.0158 6.1638
PCR-1F 0.0075 0.0376 0.2002 0.1203
PCR-3F 0.0082 0.0382 0.2151 0.2860
PCR-5F 0.0081 0.0557 0.1462 1.0298
PCR-7F 0.0128 0.0433 0.2964 1.0439
PCR-Adaptive 0.0081 0.0435 0.1865 0.4231
with transaction cost
RNW-GIC-SF 0.0056 0.0441 0.1269 NA
RNW-GIC-3F 0.0069 0.0446 0.1555 NA
RNW-GIC-5F 0.0075 0.0480 0.1564 NA
NW-GIC 0.0088 0.0429 0.2044 NA
POET 0.0059 0.0431 0.1359 NA
Ledoit-Wolf 0.0047 0.0521 0.0908 NA
Inter-RRE -0.0055 0.1166 -0.0471 NA
PCR-1F 0.0074 0.0378 0.1956 NA
PCR-3F 0.0086 0.0379 0.2263 NA
PCR-5F 0.0069 0.0560 0.1239 NA
PCR-7F 0.0116 0.0433 0.2682 NA
PCR-Adaptive 0.0074 0.0436 0.1705 NA
Table 3: Monthly Portfolio Performance of 357 Stocks, $n_{I}=240$, $n-n_{I}=108$
Estimator Maximum Sharpe Ratio
Return SD Sharpe Turnover
without transaction cost
RNW-GIC-SF 0.0046 0.0432 0.1068 0.2857
RNW-GIC-3F 0.0057 0.0440 0.1297 0.3580
RNW-GIC-5F 0.0056 0.0461 0.1209 0.4507
NW-GIC 0.0093 0.0422 0.2202 0.1333
POET 0.0058 0.0424 0.1364 0.2009
Ledoit-Wolf 0.0015 0.0534 0.0280 0.9073
Inter-RRE 0.0106 0.0453 0.2338 0.2725
PCR-1F 0.0073 0.0364 0.2004 0.1349
PCR-3F 0.0054 0.0366 0.1470 0.2798
PCR-5F 0.0056 0.0528 0.1056 1.6010
PCR-7F 0.0095 0.0506 0.1870 1.8931
PCR-Adaptive 0.0085 0.0420 0.2020 0.2608
with transaction cost
RNW-GIC-SF 0.0049 0.0430 0.1136 NA
RNW-GIC-3F 0.0059 0.0439 0.1346 NA
RNW-GIC-5F 0.0058 0.0458 0.1267 NA
NW-GIC 0.0088 0.0423 0.2082 NA
POET 0.0059 0.0424 0.1399 NA
Ledoit-Wolf 0.0015 0.0530 0.0281 NA
Inter-RRE 0.0100 0.0454 0.2203 NA
PCR-1F 0.0072 0.0366 0.1961 NA
PCR-3F 0.0057 0.0364 0.1567 NA
PCR-5F 0.0047 0.0528 0.0893 NA
PCR-7F 0.0073 0.0507 0.1436 NA
PCR-Adaptive 0.0079 0.0420 0.1872 NA
6 Conclusion
In this paper, we have introduced a general precision matrix estimator. Our estimator is model-free and does not rely on any sparsity assumptions. Under certain conditions, our estimator exhibits double descent in the estimation error. A distinguishing feature of our analysis is that it is done for finite samples, and leads to non-asymptotic bounds for the error of precision matrix estimator. Our numerical results support our theoretical findings. We have shown that the proposed estimator compares favorably with popular existing machine learning estimator alternatives. Moreover, the Sharpe ratio produced by one of our estimators on the S&P 500 data exhibits the double ascent, which to the best of our knowledge has never obtained in other empirical studies.
APPENDIX
In this appendix, we only cover the proof of Theorem 2, the rest of the proofs are relegated to Online Appendix. Before the following proof, we proceed with three inequalities that are used for rectangular matrices. Take a generic rectangular matrix $M$ and a vector $x$.
$\|Mx\|_{2}\leq\|M\|_{2}\|x\|_{2},$ (A.1)
by (11.4.4) of Bernstein (2018). Then by (11.3.15) of Bernstein (2018) where for two generic rectangular matrices $M,L$
$\|ML\|_{2}\leq\|M\|_{2}\|L\|_{2},$ (A.2)
We form the following matrix norm inequality for a generic rectangular matrix $M:m_{1}\times m_{2}$
$\|M\|_{2}\leq\|M\|_{F}$ $=$ $[\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}|M_{ij}|^{2}]^{1/2}$ (A.3)
$\leq$ $\sqrt{m_{1}m_{2}}\max_{1\leq i\leq m_{1}}\max_{1\leq j\leq m_{2}}|M_{ij}|=\sqrt{m_{1}m_{2}}\|M\|_{\infty}.$
where .$\|.\|_{F}$ is the Frobenius norm and we get the first inequality by (4.4) of Vershynin (2019).
Proof of Theorem 2.
First note that $Ey_{jt}=\mu_{j}=a_{j}^{\prime}Ef_{t}$, and $EY_{-jt}=\mu_{-j}=A_{-j}Ef_{t}$. Also see that
$y_{jt}-\mu_{j}=a_{j}^{\prime}(f_{t}-Ef_{t})+u_{jt},$ (A.4)
and
$Y_{-jt}-\mu_{-j}=A_{-j}(f_{t}-Ef_{t})+U_{-jt},$ (A.5)
(i). Note that we need to estimate the $j$ th main diagonal term for the precision matrix of asset returns. In that sense the key element is the reciprocal of that term which is defined in (12), and the population version in (OB.11). We use triangle inequality to consider the estimation error by (12)(OB.11)
$|\tilde{\tau}_{j}^{2}-\tau_{j}^{2}|\leq|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}^{2}-E(y_{jt}-\mu_{j})^{2}]|+\left|[n^{-1}\sum_{t=1}^{n}y_{jt}Y_{-jt}^{\prime}]\tilde{\alpha}_{j}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}\alpha_{j}^{*}\right|.$ (A.6)
Next add and subtract to the second right side term $E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\tilde{\alpha}_{j}$, by iid nature of data,
$|\tilde{\tau}_{j}^{2}-\tau_{j}^{2}|$$\leq$ $|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}^{2}-E(y_{jt}-\mu_{j})^{2}]|+\left|n^{-1}\sum_{t=1}^{n}[y_{jt}Y_{-jt}^{\prime}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}\tilde{\alpha}_{j}\right|$ (A.7)
$+$ $\left|E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}](\tilde{\alpha}_{j}-\alpha_{j}^{*})\right|.$
Take uniform bounds over $j=1,\cdots,p$ and use Cauchy-Schwartz inequality for the second and third right side terms on (A.7)
$\max_{1\leq j\leq p}|\tilde{\tau}_{j}^{2}-\tau_{j}^{2}|$$\leq$ $\max_{1\leq j\leq p}|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}^{2}-E(y_{jt}-\mu_{j})^{2}]|$ (A.8)
$+$ $\max_{1\leq j\leq p}\|n^{-1}\sum_{t=1}^{n}[y_{jt}Y_{-jt}^{\prime}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}\|_{2}$
$+$ $\max_{1\leq j\leq p}\|E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}.$
Each right side term in (A.8) will be analyzed as a step in the proof.
Step 1. Start with the first right side term on (A.8), and by (2)
$y_{jt}^{2}=a_{j}^{\prime}f_{t}f_{t}^{\prime}a_{j}+2a_{j}^{\prime}f_{t}u_{jt}+u_{jt}^{2},$
and by Assumption 1(iv), and using (A.4)
$E(y_{jt}-\mu_{j})^{2}=a_{j}^{\prime}\Sigma_{f}a_{j}+E(u_{jt}^{2}).$
Use triangle inequality by the last two equalities, and noting $\Sigma_{f}:=E(f_{t}-Ef_{t})(f_{t}-Ef_{t})^{\prime}$
$\max_{1\leq j\leq p}|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}^{2}-E(y_{jt}-\mu_{j})^{2}]|$$\leq$ $\max_{1\leq j\leq p}|a_{j}^{\prime}[\frac{1}{n}\sum_{t=1}^{n}(f_{t}f_{t}^{\prime}-\Sigma_{f})]a_{j}|$ (A.9)
$+$ $\max_{1\leq j\leq p}|\frac{1}{n}\sum_{t=1}^{n}[u_{jt}^{2}-E(u_{jt}^{2})]|$
$+$ $\max_{1\leq j\leq p}2|a_{j}^{\prime}[n^{-1}\sum_{t=1}^{n}f_{t}u_{jt}]|.$
Take the first right side term in (A.9) above, and use Cauchy-Schwarz inequality first and then (A.1)
$\max_{1\leq j\leq p}|a_{j}^{\prime}[\frac{1}{n}\sum_{t=1}^{n}(f_{t}f_{t}^{\prime}-\Sigma_{f})]a_{j}|\leq\left[\max_{1\leq j\leq p}\|a_{j}\|_{2}^{2}\right]\|\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]\|_{2}.$ (A.10)
We consider the second term on the right side of (A.10), with $F:n\times K$ matrix, and $\|\Sigma_{f}\|_{2}\leq C<\infty$ by Assumption 3(ii)
$\|\frac{F^{\prime}F}{n}-\Sigma_{f}\|_{2}\leq C[\sqrt{\frac{K}{n}}+\sqrt{\frac{logn}{n}}]\|\Sigma_{f}\|_{2},$ (A.11)
by Lemma 23 of Bing et al. (2021) since rows of $F$ are independent subgaussian vectors, with probability at least $1-2/n^{c}$ via Assumption 1(iii), Assumption 4(ii), and note that $K<n$. By Assumption 4(iii)
$\max_{1\leq j\leq p}\|a_{j}\|_{2}^{2}\leq CK.$ (A.12)
Use (A.11)-(A.12) in (A.10) and Assumption 3 to have
$\max_{1\leq j\leq p}|a_{j}^{\prime}(\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]a_{j}|=O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\sqrt{\frac{logn}{n}})\right).$ (A.13)
Then in (A.9) consider the second right side term
$\max_{1\leq j\leq p}|\frac{1}{n}\sum_{t=1}^{n}[u_{jt}^{2}-E(u_{jt}^{2})]|=O_{p}(\sqrt{\frac{logp}{n}}).$ (A.14)
We now show how we obtain (A.14). Note that by Corollary 2.8.3 of Vershynin (2019), for $t_{1}>0$, and the union bound
$P\left[\max_{1\leq j_{1}\leq p}\max_{1\leq j_{2}\leq p}|\frac{1}{n}\sum_{t=1}^{n}[u_{j_{1},t}u_{j_{2},t}-E(u_{j_{1},t}u_{j_{2},t})]|>t_{1}\right]\leq p^{2}exp\left(-cmin(\frac{t_{1}^{2}}{C_{\psi}^{2}},\frac{t_{1}}{C_{\psi}})n\right),$
where we use $u_{j_{1},t}$ and $u_{j_{2},t}$ are both subgaussian random variables, so $u_{j_{1},t}u_{j_{2},t}$ are subexponential random variable (independent across $t=1,\cdots,n$) by Lemma 2.7.7 of Vershynin (2019) and exercise 2.7.10 of Vershynin (2019) shows centered subexponential are also subexponential random variables, and we have
$\max_{1\leq j_{1}\leq p}\max_{1\leq j_{2}\leq p}\|u_{j_{1},t}u_{j_{2},t}\|_{\psi}\leq C_{\psi}<\infty$
where $C_{\psi}$ are the upper bound constant for the subexponential norm in Definition 2.7.5 of Vershynin (2019). Next $2p^{2}exp\left(-cmin(\frac{t_{1}}{C_{\psi}^{2}},\frac{t_{1}}{C_{\psi}})n\right)=exp\left(log(2p^{2})-c\frac{t_{1}^{2}n}{C_{\psi}^{2}}\right)$, with $t_{1}:=C\frac{\sqrt{logp}}{\sqrt{n}}$ and with sufficiently large $n$ (i.e.$\sqrt{n}\geq\frac{C}{C_{\psi}}(logp)^{1/2}$). Set $c_{\psi}:=c\frac{C^{2}}{C_{\psi}^{2}}-2>0$ to have $exp[log2+2logp-c\frac{C^{2}}{C_{\psi}^{2}}logp]=\frac{2}{p^{c_{\psi}}}.$ Then
$P\left(\max_{1\leq j_{1}\leq p}\max_{1\leq j_{2}\leq p}|\frac{1}{n}\sum_{t=1}^{n}u_{j_{1},t}u_{j_{2},t}-Eu_{j_{1},t}u_{j_{2},t}|>C\frac{\sqrt{logp}}{\sqrt{n}}\right)\leq\frac{2}{p^{c_{\psi}}}.$
The asymptotics in (A.14) are obtained after taking $n\to\infty,p\to\infty$. Consider the third right side term in (A.9). Lemma 2.7.7 and Corollary 2.8.3 of Vershynin (2019) together with Assumptions 1, 4, and Holder’s inequality give
$\max_{1\leq j\leq p}|a_{j}^{\prime}(\frac{1}{n}\sum_{t=1}^{n}f_{t}u_{jt})|$$\leq$ $\max_{1\leq j\leq p}\|a_{j}\|_{1}\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}f_{t}u_{jt}\|_{\infty}$ (A.15)
$=$ $O_{p}(K\sqrt{\frac{log(Kp)}{n}})=O_{p}(\frac{K\sqrt{logp}}{\sqrt{n}}),$
where the last equality uses $p-1>K$. By (A.13)-(A.15) we have the first right side term in (A.8)
$\max_{1\leq j\leq p}|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}^{2}-E(y_{jt}-\mu_{j})^{2}]|=O_{p}\left(\max(\frac{K^{3/2}}{n^{1/2}},K\sqrt{\frac{\max(logp,logn)}{n}})\right).$ (A.16)
Step 2. We consider the second right side term in (A.8). By (2)(3) and Assumption 1(iv)
$y_{jt}Y_{-jt}^{\prime}=a_{j}^{\prime}f_{t}f_{t}^{\prime}A_{-j}^{\prime}+a_{j}^{\prime}f_{t}U_{-jt}^{\prime}+u_{jt}f_{t}^{\prime}A_{-j}^{\prime}+u_{jt}U_{-jt}^{\prime},$
and by (A.4) (A.5)
$E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}=a_{j}^{\prime}\Sigma_{f}A_{-j}^{\prime}.$
Use these last two equations
$\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}Y_{-jt}^{\prime}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\leq\max_{1\leq j\leq p}\|a_{j}^{\prime}\left(\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]\right)A_{-j}^{\prime}\|_{2}$ (A.17)
$+$ $\max_{1\leq j\leq p}\|(\frac{1}{n}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime})A_{-j}^{\prime}\|_{2}+\max_{1\leq j\leq p}\|a_{j}^{\prime}(\frac{1}{n}\sum_{t=1}^{n}f_{t}U_{-jt}^{\prime})\|_{2}+\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}u_{jt}U_{-jt}^{\prime}\|_{2}.$
Consider the first term on the right side of (A.17) by (A.1)(A.2)
$\max_{1\leq j\leq p}\|a_{j}^{\prime}(\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]A_{-j}^{\prime}\|_{2}\leq\max_{1\leq j\leq p}\|a_{j}\|_{2}\max_{1\leq j\leq p}\left\|\left(\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]\right)A_{-j}^{\prime}\right\|_{2}$ (A.18)
$\leq$ $(\max_{1\leq j\leq p}\|a_{j}\|_{2})\|\frac{1}{n}\sum_{t=1}^{n}[f_{t}f_{t}^{\prime}-\Sigma_{f}]\|_{2}\times(\max_{1\leq j\leq p}\|A_{-j}^{\prime}\|_{2})$
$=$ $O(K^{1/2})O_{p}(\max(\frac{K^{1/2}}{n^{1/2}},\frac{\sqrt{logn}}{n^{1/2}}))O(d_{1n})=O_{p}(max(\frac{Kd_{1n}}{n^{1/2}},\frac{d_{1n}\sqrt{Klogn}}{n^{1/2}})),$
and the rates are by (A.11)(A.12) and Assumption 3. Then consider the second term on the right side of (A.17)
$\max_{1\leq j\leq p}\|(\frac{1}{n}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime})A_{-j}^{\prime}\|_{2}$$\leq$ $\max_{1\leq j\leq p}\|(\frac{1}{n}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime})\|_{2}\max_{1\leq j\leq p}\|A_{-j}^{\prime}\|_{2}$ (A.19)
$\leq$ $\sqrt{K}\max_{1\leq j\leq p}\|(\frac{1}{n}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime})\|_{\infty}\max_{1\leq j\leq p}\|A_{-j}^{\prime}\|_{2}$
$=$ $O_{p}(\sqrt{K}\frac{\sqrt{logp}}{\sqrt{n}})O(d_{1n}),$
where we use (A.1) for the first inequality since $u_{jt}f_{t}^{\prime}:1\times K$ vector and $A_{-j}^{\prime}:K\times p-1$ matrix, and the second inequality is by simple vector norm inequality tying $l_{2},l_{\infty}$ vector norms, and the rates are by Assumption 3(iii) $\max_{1\leq j\leq p}\|A_{-j}^{\prime}\|_{2}=O(d_{1n})$ and Lemma 2.7.7 and Corollary 2.8.3. of Vershynin (2019) which provides $\|n^{-1}\sum_{t=1}^{n}u_{jt}f_{t}^{\prime}\|_{\infty}=O_{p}(\sqrt{logp/n})$. Now third term on the right side of (A.17)
$\max_{1\leq j\leq p}\|a_{j}^{\prime}(\frac{1}{n}\sum_{t=1}^{n}f_{t}U_{-jt}^{\prime})\|_{2}$$\leq$ $\max_{1\leq j\leq p}\|a_{j}\|_{2}\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}f_{t}U_{-jt}^{\prime}\|_{2}$ (A.20)
$\leq$ $\max_{1\leq j\leq p}\|a_{j}\|_{2}\sqrt{Kp}\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}f_{t}U_{-jt}^{\prime}\|_{\infty}$
$=$ $O(\sqrt{K})O_{p}(\frac{\sqrt{Kp(logp)}}{\sqrt{n}})=O_{p}(K\sqrt{p}\frac{\sqrt{logp}}{\sqrt{n}}),$
where we use for the first inequality (A.1) and then for the second inequality we use (A.3) and for the rates we use Assumption 5(iii), and Lemma 2.7.7 and Corollary 2.8.3 of Vershynin (2019) with Assumption 4 since $f_{t}U_{-jt}^{\prime}$ is $K\times p-1$ matrix. Then for the fourth term on the right side of (A.17) via Corollary 2.8.3 of Vershynin (2019) and Assumption 4 and simple vector norm inequality
$\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}u_{jt}U_{-jt}^{\prime}\|_{2}\leq\sqrt{p}\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}u_{jt}U_{-jt}^{\prime}\|_{\infty}=O_{p}(\sqrt{p}\frac{\sqrt{logp}}{\sqrt{n}}).$ (A.21)
Relying on (A.18)(A.19)(A.20)(A.21), and keeping in mind that (A.20) has the largest rate by Assumptions 3 and 5(i), and we have the following for the second right side term in (A.8)
$\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}Y_{-jt}^{\prime}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}=O_{p}(K\sqrt{p}\frac{\sqrt{logp}}{\sqrt{n}}).$ (A.22)
We find it useful to recall the rate definition
$r_{w1}:=\max\left(\sqrt{\frac{logn+\bar{r}}{n\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}},\sqrt{\frac{K\bar{\Psi}}{\bar{\xi}\bar{\eta}\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}\bar{\eta}}}\right),$ (A.23)
The proof of Theorem 1-Step 2d together with $r_{n}/\delta_{n}\to 1$, gives
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}\|_{2}=O_{p}(r_{w1}).$ (A.24)
Combine (A.22)(A.24) for the second right side term in (A.8)
$\max_{1\leq j\leq p}\|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}Y_{-jt}^{\prime}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}\|_{2}$ (A.25)
$=$ $O_{p}((K\sqrt{plogp/n})r_{w1})$
Step 3. Now consider the third term on the right side of (A.8). The key term is $E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]=a_{j}\Sigma_{f}A_{-j}^{\prime},$ where we use Assumption 1 and (2)(3). Since we have for rectangular matrices (A.1)(A.2)
$\max_{1\leq j\leq p}\|E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}\|_{2}=\max_{1\leq j\leq p}\|a_{j}^{\prime}\Sigma_{f}A_{-j}^{\prime}\|_{2}=d_{2n},$ (A.26)
By (A.26), the third term on the right side of (A.8) is by Theorem 1
$\max_{1\leq j\leq p}\|E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}\|_{2}\max_{1\leq j\leq p}\|\hat{\alpha}_{j}-\alpha_{j}^{*}\|_{2}=O_{p}(d_{2n}r_{w2}),$ (A.27)
with $r_{w2}:=\max(r_{w1},\sqrt{\frac{K}{\bar{\xi}}})$.
Step 4. We combine (A.16)(A.25)(A.27) with Assumption 5(i) on the right side of (A.8) to have the desired result.
Q.E.D.
(ii). By Assumption 5(ii)-(iii) we obtain the consistency. Q.E.D.
(iii). By (B.11) $\tau_{j}^{2}:=E[(y_{jt}-\mu_{j})^{2}-E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\alpha_{j}^{*}.$ Then by Cauchy-Schwarz inequality
$\max_{1\leq j\leq p}E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\alpha_{j}^{*}$$\leq$ $\left[\max_{1\leq j\leq p}\|E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]\|_{2}\right]\left[\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}\right]$ (A.28)
$\leq$ $\frac{CK^{1/2}d_{2n}}{\sqrt{\bar{\xi}}}\to 0,$
where we use Step 3 of Theorem 2 together with Step 3 of Theorem 1 proof, and (A.26), and Assumption 5(iii). Next by using (2), Assumptions 1-3
$\min_{1\leq j\leq p}E(y_{jt}-\mu_{j})^{2}$$\geq$ $min_{1\leq j\leq p}a_{j}^{\prime}\Sigma_{f}a_{j}+\min_{1\leq j\leq p}E(u_{jt}^{2})$
$\geq$ $Eigmin(\Sigma_{f})\min_{1\leq j\leq p}\|a_{j}\|_{2}^{2}+\min_{1\leq j\leq p}\sigma_{j}^{2}\geq c>0.$
The last two results provide the desired result. Q.E.D.
Appendix A ONLINE APPENDIX
The Online Appendix is organized into three sections. Appendix A contains the main proofs from the text (excluding Theorem 2, which appears in the main text’s appendix). Appendix B provides supporting results related to matrix algebra, including inverse matrix identities and two tail inequalities. Online Appendix C presents additional simulations and empirical results.
ONLINE APPENDIX A
Proof of Theorem 1:
Using (4) and (11) we get
$\tilde{\alpha}_{j}-\alpha_{j}^{*}=\hat{B}(Y_{-j}\hat{B})^{+}Fa_{j}+\hat{B}(Y_{-j}\hat{B})^{+}u_{j}-\alpha_{j}^{*}.$
Then we have the following inequality
$\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}\leq 3\|\hat{B}(Y_{-j}\hat{B})^{+}u_{j}\|_{2}^{2}+3\|\hat{B}(Y_{-j}\hat{B})^{+}Fa_{j}\|_{2}^{2}+3\|\alpha_{j}^{*}\|_{2}^{2}.$ (OA.1)
The proof contains a detailed transformation of results that can be obtained in linear regression framework to precision matrix estimation needed here (to make connection with linear regression clearer and overall exposition easier to follow, we often refer to Lemma 17 and related results of Bing et al. (2021). Since we ultimately aim to use our results for high dimensional outcomes, we need them to hold uniformly over $j=1,\dots,p,$. It is important to note that this is in striking contrast with pointwise results (as are those from, e.g., Bing et al. (2021) that would be applicable to only one outcome. Each of the right side terms in (OA.1) will be analyzed in each of Steps 1-3.
Step 1. Start with the first right side term in (OA.1)
$\|\hat{B}(Y_{-j}\hat{B})^{+}u_{j}\|_{2}^{2}=u_{j}^{\prime}M_{j}u_{j},$ (OA.2)
with $M_{j}:n\times n$ matrix defined as
$M_{j}:=(Y_{-j}\hat{B})^{+^{\prime}}\hat{B}^{\prime}\hat{B}(Y_{-j}\hat{B})^{+}.$ (OA.3)
To simplify the proof, , we consider the following trace term: $tr(M_{j})=tr[(Y_{-j}\hat{B})^{+^{\prime}}\hat{B}^{\prime}\hat{B}(Y_{-j}\hat{B})^{+}].$ First of all
$tr[M_{j}]$$\leq$ $rank(M_{j})\|M_{j}\|_{2}=rank(Y_{-j}\hat{B})^{+}\|M_{j}\|_{2}$ (OA.4)
$=$ $rank(Y_{-j}\hat{B})\|M_{j}\|_{2}=rank(Y_{-j}P_{\hat{B}})\|M_{j}\|_{2},=\hat{r}_{j}\|M_{j}\|_{2}$
where the first inequality is due to $M_{j}$ being positive semidefinite, and $n\times n$ matrix and Exercise 7.49 of Abadir and Magnus (2005), and the first equality is by Exercise 8.27a and 4.24a of Abadir and Magnus (2005), and the second equality is by Exercise 10.28 of Abadir and Magnus (2005), and the third equality is again by Exercises 4.13d, 4.15b, 4.25b, and 10.28 of Abadir and Magnus (2005) with $P_{\hat{B}}=\hat{B}\hat{B}^{+}$ in p.8 of Bing et al. (2021), and the last equality is just the definition of the complexity $\hat{r}_{j}$. Now we consider, as proved in (44) of Bing et al. (2021)
$\|M_{j}\|_{2}\leq(n\hat{\eta}_{j})^{-1},$ (OA.5)
where $\hat{\eta}_{j}:=\hat{\sigma}_{\hat{r}_{j}}^{2}(Y_{-j}P_{\hat{B}})/n$ which is the $\hat{r}_{j}$ th largest singular value, squared, of the matrix $Y_{-j}P_{\hat{B}}$, and divided by $n$.
Define $C_{p}>0$ a positive constant. Then by (OA.4)(OA.5) and $\bar{\eta}$, $\bar{r}$ definitions
$\gamma_{e}^{2}\sigma_{j}^{2}[2C_{p}\|M_{j}\|_{2}logn+tr(M_{j})]\leq\gamma_{e}^{2}\sigma^{2}[\frac{2C_{p}logn}{n\bar{\eta}}+\frac{\bar{r}}{n\bar{\eta}}].$ (OA.6)
Next, define the following events
$E_{1j}:=\{u_{j}^{\prime}Mu_{j}\leq 2\gamma_{e}^{2}\sigma_{j}^{2}[2C_{p}\|M_{j}\|_{2}logn+tr(M_{j})]\},$ (OA.7)
and
$\bar{E}_{1j}:=\{u_{j}^{\prime}Mu_{j}\leq 2\gamma_{e}^{2}\sigma^{2}[\frac{2C_{p}logn}{n\bar{\eta}}+\frac{\bar{r}}{n\bar{\eta}}]\}.$ (OA.8)
with
$\bar{E}_{1}:=\{\max_{1\leq j\leq p}u_{j}^{\prime}Mu_{j}\leq 2\gamma_{e}^{2}\sigma^{2}[\frac{2C_{p}logn}{n\bar{\eta}}+\frac{\bar{r}}{n\bar{\eta}}]\}.$ (OA.9)
By (OA.6)
$E_{1j}\subseteq\bar{E}_{1j}.$ (OA.10)
Clearly, since the upper bounds of the events $\bar{E}_{1j},\bar{E}_{1}$ are the same
$\cap_{j=1}^{p}\bar{E}_{1j}=\bar{E}_{1}.$ (OA.11)
By (OA.10)(OA.11)
$\cap_{j=1}^{p}E_{1j}\subseteq\bar{E_{1}}.$ (OA.12)
See that by $E_{1j}^{c}$ as the complement of $E_{1j}$
$P(\cap_{j=1}^{p}E_{1j})+P(\cup_{j=1}^{p}E_{1j}^{c})=1.$ (OA.13)
Note that we have, by (OA.12)
$P(\bar{E}_{1})\geq P(\cap_{j=1}^{p}E_{1j})$ (OA.14)
so by (OA.13)(OA.14) and
$P(\bar{E}_{1})$$\geq$ $1-P(\cup_{j=1}^{p}E_{1j}^{c})$ (OA.15)
$\geq$ $1-\sum_{j=1}^{p}P(E_{1j}^{c}).$
Now we calculate the sum of the upper bound probability of $E_{1j}^{c}$, across $j$. Note that $u_{j}$ is a subgaussian vector, and independent of $Y_{-j}$, and $\hat{B}$ by Assumption 1 and the definition of $\hat{B}$, with $M_{j}$ being a symmetric positive semi-definite matrix. So $M_{j}$ is independent of $u_{j}$. Then Lemma OB.1 (with $t=C_{p}logn$, $H=M_{j},\zeta=u_{j}$, conditional on $M_{j}$) allows us to write
$P(E_{1j}^{c}|M_{j})=P\left(u_{j}^{\prime}Mu_{j}>2\gamma_{e}^{2}\sigma_{j}^{2}[2C_{p}\|M_{j}\|_{2}logn+tr(M_{j})]|M_{j}\right)\leq\frac{1}{n^{C_{p}}}.$ (OA.16)
Next, use the union bound and by Assumption 3, using $logp\leq(\frac{C_{p}}{2})logn$
$\sum_{j=1}^{p}P(E_{1j}^{c})$ $=$ $\sum_{j=1}^{p}P(E_{1j}^{c}|M_{j})P(M_{j})\leq\sum_{j=1}^{p}P(E_{1j}^{c}|M_{j})$ (OA.17)
$\leq$ $exp(logp-C_{p}logn)\leq exp(-\frac{C_{p}}{2}logn)=\frac{1}{n^{C_{p}/2}}.$
Use (OA.15)(OA.17) in (OA.2)(OA.9)
$P\left(\max_{1\leq j\leq p}\|\hat{B}(Y_{-j}\hat{B})^{+}u_{j}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}[\frac{2C_{p}logn}{n\bar{\eta}}+\frac{\bar{r}}{n\bar{\eta}}]\right)\geq 1-\frac{1}{n^{C_{p}/2}}.$ (OA.18)
Step 2. Now consider the second term on the right side of (OA.1). Keeping in mind that, $A_{-j}^{\prime}(A_{-j}^{+})^{\prime}=I_{K}$ and recalling on (5), while setting $M_{\hat{B}}:=I_{p-1}-P_{\hat{B}}$, we further have
$\hat{B}(Y_{-j}\hat{B})^{+}F=\hat{B}(Y_{-j}\hat{B})^{+}FA_{-j}^{\prime}A_{-j}^{+^{\prime}}=\hat{B}(Y_{-j}\hat{B})^{+}(Y_{-j}-U_{-j})A_{-j}^{+^{\prime}}$
$=$ $\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}A_{-j}^{+^{\prime}}+\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}M_{\hat{B}}A_{-j}^{+^{\prime}}-\hat{B}(Y_{-j}\hat{B})^{+}U_{-j}A_{j}^{+^{\prime}}$
Use simple inequality $(a+b+c)^{2}\leq 3a^{2}+3b^{2}+3c^{2}$,
$\|\hat{B}(Y_{-j}\hat{B})^{+}Fa_{j}\|_{2}^{2}\leq 3\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}$ (OA.19)
$+$ $3\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}(I_{p-1}-P_{\hat{B}})A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}+3\|\hat{B}(Y_{-j}\hat{B})^{+}U_{-j}A_{j}^{+^{\prime}}a_{j}\|_{2}^{2}.$
In Steps 2a, 2b, and 2c below, we separetely analyze each of the three terms on the right side of (OA.19).
Step 2a. So
$\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}\|_{2}^{2}\|A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}.$ (OA.20)
Note that we can write the first term on the right side of (OA.20) as
$\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}\|_{2}=\|\hat{B}(\hat{B}^{\prime}Y_{-j}^{\prime}Y_{-j}\hat{B})^{+}\hat{B}^{\prime}Y_{-j}^{\prime}Y_{-j}P_{\hat{B}}\|_{2},$ (OA.21)
where we use Moore-Penrose inverse rule (134) in Bunea et al. (2022) which is for a rectangular matrix $S$, we have $(S^{\prime}S)^{+}S^{\prime}=S^{+}$, which we input $S=Y_{-j}\hat{B}$ in our case above. To simplify the expressions see that by the proof of Lemma 15 of Bing et al. (2021) $P_{\hat{B}}=V_{1}V_{1}^{\prime},$ where $\hat{B}=V_{1}DV_{2}^{\prime}$ is a singular value decomposition. $V_{1},V_{2}$ are orthonormal matrices and with dimensions of $p-1\times r_{0},q\times r_{0}$ respectively, and $D$ is a $r_{0}$ dimensional diagonal matrix, and the rank of $(\hat{B})=r_{0}$. Rewrite right side of (OA.21) as
$\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}\|_{2}=\|V_{1}DV_{2}^{\prime}(V_{2}DV_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}DV_{2}^{\prime})^{+}V_{2}DV_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}V_{1}^{\prime}\|_{2}.$ (OA.22)
Then in (OA.22) right side see that
$(V_{2}DV_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}DV_{2}^{\prime})^{+}=V_{2}D^{-1}(V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1})^{+}D^{-1}V_{2}^{\prime},$ (OA.23)
which is shown in the proof of Lemma 15 of Bing et al. (2021). Then use (OA.23) in (OA.22), $V_{2}^{\prime}V_{2}=I_{r_{0}}$
$V_{1}DV_{2}^{\prime}(V_{2}DV_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}DV_{2}^{\prime})^{+}V_{2}DV_{1}^{\prime}$ $=$ $V_{1}DV_{2}^{\prime}V_{2}D^{-1}[V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}]^{+}D^{-1}V_{2}^{\prime}V_{2}DV_{1}^{\prime}$
$=$ $V_{1}[V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}]^{+}V_{1}^{\prime}$
Use the last expression in (OA.22) right side
$\|V_{1}DV_{2}^{\prime}(V_{2}DV_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}DV_{2}^{\prime})^{+}V_{2}DV_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}V_{1}^{\prime}\|_{2}=\|V_{1}[V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}]^{+}V_{1}^{\prime}Y_{-j}Y_{-j}^{\prime}V_{1}V_{1}^{\prime}\|_{2}.$
Then clearly since $V_{1}:p-1\times r_{0}$ matrix, and Exercise 10.26 of Abadir and Magnus (2005) shows $[V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}]^{+}V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}$ is a symmetric idempotent matrix and all idempotent matrices eigenvalues are either 0 or 1, and since $V_{1}^{\prime}V_{1}=I_{r_{0}}$
$\|V_{1}[V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}]^{+}V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}V_{1}^{\prime}\|_{2}\leq\|V_{1}\|_{2}\|[V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}]^{+}V_{1}^{\prime}Y_{-j}^{\prime}Y_{-j}V_{1}\|_{2}\|V_{1}^{\prime}\|_{2}\leq 1.$
So use this last inequality in (OA.22)
$\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}\|_{2}\leq 1,$ (OA.24)
Next part of the proof simplifies certain expressions in Lemma 17 of Bing et al. (2021) and ties that to signal-to-noise ratio. Now consider the following term which is common in almost all terms on upper bounds of our estimator, and will link to signal to noise ratio, which is the second term on the right side of (OA.20), $\|A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}=a_{j}^{\prime}A_{-j}^{+}A_{-j}^{+^{\prime}}a_{j}=a_{j}^{\prime}(A_{-j}^{\prime}A_{-j})^{-1}a_{j},$ since $A_{-j}^{+}=(A_{-j}^{\prime}A_{-j})^{-1}A_{-j}^{{}^{\prime}}$ with $p-1>K$. Next
$a_{j}^{\prime}A_{-j}^{+}(A_{-j}^{+})^{\prime}a_{j}$ $=$ $a_{j}^{\prime}\Sigma_{f}^{1/2}(\Sigma_{f}^{1/2}A_{-j}^{\prime}A_{-j}\Sigma_{f}^{1/2})^{-1}\Sigma_{f}^{1/2}a_{j}$ (OA.25)
$\leq$ $\frac{a_{j}^{\prime}\Sigma_{f}a_{j}}{\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})},$
where we use first Rayleigh inequality on p.182 of Abadir and Magnus (2005) and then Theorem 2.6.3b of Horn and Johnson (2013), where $\lambda_{K}(.)$ is the $K$ th largest eigenvalue for the matrix $A_{-j}\Sigma_{f}A_{-j}^{\prime}:p-1\times p-1$. By (6)(7), we try to rewrite (OA.25) right-side.
$\lambda_{K}(A_{-j}^{\prime}\Sigma_{f}A_{-j})$ $=$ $\left[\frac{\lambda_{K}(A_{-j}^{\prime}\Sigma_{f}A_{-j})}{\|\Sigma_{U,-j}\|_{2}}\right]\|\Sigma_{U,-j}\|_{2}$ (OA.26)
$=$ $\xi_{j}\|\Sigma_{U,-j}\|_{2}\geq\bar{\xi}\|\Sigma_{U,-j}\|_{2}$
$\geq$ $\bar{\xi}\delta_{n},$
where we use $\delta_{n}:=\min_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}\geq c>0$ by Assumption 2(ii), where $\delta_{n}$ is non-decreasing in $n$. Then by $\max_{1\leq j\leq p}\max_{1\leq k\leq K}|a_{jk}|\leq C<\infty$, with $a_{jk}$ as the $k$ th element of row $a_{j}$ in $A$, which is Assumption 4(iii)
$a_{j}^{\prime}\Sigma_{f}a_{j}\leq\|a_{j}\|_{2}^{2}Eigmax(\Sigma_{f})\leq CK,$ (OA.27)
by $Eigmax(\Sigma_{f})\leq C<\infty$ Assumption 3. So the first term on the right side of (OA.25) can be bounded as as
$\max_{1\leq j\leq p}\|A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}=\max_{1\leq j\leq p}\left[\frac{(a_{j}^{\prime}\Sigma_{f}a_{j})}{\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})}\right]\leq\frac{CK}{\bar{\xi}\delta_{n}}.$ (OA.28)
Clearly by (OA.24)(OA.28) in (OA.20)
$\max_{1\leq j\leq p}\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}P_{\hat{B}}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq\frac{CK}{\bar{\xi}\delta_{n}}.$ (OA.29)
Step 2b. Take the second term on the right side of (OA.19)
$\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}(I_{p-1}-P_{\hat{B}})A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq\|\hat{B}(Y_{-j}\hat{B})^{+}\|_{2}^{2}\|Y_{-j}(I_{p-1}-P_{\hat{B}})A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}.$ (OA.30)
See that the first term on the right side in (OA.30) by $\|S_{b}\|_{2}^{2}=\|S_{b}^{\prime}S_{b}\|_{2}$ for a rectangular matrix $S_{b}:=\hat{B}(Y_{-j}\hat{B})^{+}$ and $M_{j}$ definition in (OA.3)
$\|\hat{B}(Y_{-j}\hat{B})^{+}\|_{2}^{2}=\|M_{j}\|_{2}\leq\frac{1}{n\hat{\eta}_{j}},$ (OA.31)
where the last inequality is by (OA.5). Then by definition of $\hat{\Psi}_{j}$ in (17)
$\|Y_{-j}(I_{p-1}-P_{\hat{B}})\|_{2}^{2}=\sigma_{1}^{2}[(Y_{-j}(I_{p-1}-P_{\hat{B}})]=n\hat{\Psi}_{j}.$ (OA.32)
Now consider (OA.28)(OA.31)(OA.32) on the second term-right side of (OA.19), with
$\bar{\Psi}:=\max_{1\leq j\leq p}\hat{\Psi}_{j},\bar{\eta}:=\min_{1\leq j\leq p}\hat{\eta}_{j},$
with $\hat{\eta}_{j}$ defined in (16)
$\max_{1\leq j\leq p}\|\hat{B}(Y_{-j}\hat{B})^{+}Y_{-j}(I_{p-1}-P_{\hat{B}})A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq C\left(\frac{\bar{\Psi}}{\bar{\eta}}\right)\left(\frac{K}{\bar{\xi}\delta_{n}}\right).$ (OA.33)
Step 2c. Consider the third term on the right side of (OA.19)
$\|\hat{B}(Y_{-j}\hat{B})^{+}U_{-j}A_{j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq\|\hat{B}(Y_{-j}\hat{B})^{+}\|_{2}^{2}\|U_{-j}A_{j}^{+^{\prime}}a_{j}\|_{2}^{2},$ (OA.34)
where we analyze the following on the right side of (OA.34): $\|U_{-j}A_{j}^{+^{\prime}}a_{j}\|_{2}^{2}=a_{j}^{\prime}A_{-j}^{+}U_{-j}U_{-j}^{\prime}A_{-j}^{+^{\prime}}a_{j}.$ Define $\gamma_{e,j}^{2}=\gamma_{w}^{2}a_{j}^{\prime}A_{j}^{+}\Sigma_{U,-j}A_{j}^{+^{\prime}}a_{j}$. Now we simplify $\gamma_{e,j}^{2}$ term. See that, uniformly over $j$, by Assumption 2 and (OA.28)
$\max_{1\leq j\leq p}\gamma_{w}^{2}a_{j}^{\prime}A_{j}^{+}\Sigma_{U,-j}A_{j}^{+^{\prime}}a_{j}$$\leq$ $[max_{1\leq j\leq p}Eigmax(\Sigma_{U,-j})][\max_{1\leq j\leq p}\gamma_{w}^{2}a_{j}^{\prime}A_{j}^{+}A_{j}^{+^{\prime}}a_{j}]$ (OA.35)
$\leq$ $\frac{\gamma_{w}^{2}Cr_{n}K}{\bar{\xi}\delta_{n}}.$
Then define the events
$E_{2j}:=\{\|U_{-j}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq(n+2C_{p}logn+2\sqrt{C_{p}nlogn})\gamma_{e,j}^{2}\}$ (OA.36)
and
$\bar{E}_{2j}:=\{\|U_{-j}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq(n+2C_{p}logn+2\sqrt{C_{p}nlogn})\frac{C\gamma_{w}^{2}r_{n}K}{\bar{\xi}\delta_{n}}\}.$ (OA.37)
By (OA.35) and the explanations above that equation, with (OA.36)(OA.37)
$E_{2j}\subseteq\bar{E}_{2j}.$ (OA.38)
Next, define
$\bar{E}_{2}:=\{\max_{1\leq j\leq p}\|U_{-j}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq(n+2C_{p}logn+2\sqrt{C_{p}nlogn})\frac{C\gamma_{w}^{2}r_{n}K}{\bar{\xi}\delta_{n}}\}.$ (OA.39)
See that by definitions in (OA.37)(OA.39)
$\cap_{j=1}^{p}\bar{E}_{2j}=\bar{E}_{2}.$ (OA.40)
Then by (OA.38)(OA.40)
$\cap_{j=1}^{p}E_{2j}\subseteq\bar{E}_{2}.$ (OA.41)
Then by (OA.41)
$P(\bar{E}_{2})\geq P(\cap_{j=1}^{p}E_{2j})=1-P(\cup E_{2j}^{c})\geq 1-\sum_{j=1}^{p}P(E_{2j}^{c}).$ (OA.42)
Since $U_{-j}A_{-j}^{+^{\prime}}a_{j}$ is $\gamma_{w}\sqrt{a_{j}^{\prime}A_{-j}^{+}\Sigma_{U,-j}A_{-j}^{+^{\prime}}a_{j}}$ subgaussian vector for each $j=1,\cdots,p$ via Assumption 4, apply Lemma OB.1 with $\epsilon_{j}:=U_{-j}A_{j}^{+^{\prime}}a_{j}$ and $H:=I_{n}$ and $\gamma_{e,j}^{2}=\gamma_{w}^{2}a_{j}^{\prime}A_{j}^{+}\Sigma_{U,-j}A_{j}^{+^{\prime}}a_{j}$ with $t=C_{p}logn$
$P[\|U_{-j}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}>(n+2C_{p}logn+2\sqrt{C_{p}nlogn})\gamma_{e,j}^{2}]\leq 1/n^{C_{p}}.$ (OA.43)
Then take the union bound and with Assumption 3, $logp\leq\frac{C_{p}}{2}logn$
$\sum_{j=1}^{p}P(E_{2j}^{c})\leq exp(logp-C_{p}logn)\leq exp(\frac{-C_{p}}{2}logn)=\frac{1}{n^{C_{p}/2}}.$ (OA.44)
Next use (OA.42)(OA.44)(OA.39)
$P(\max_{1\leq j\leq p}\|U_{-j}A_{-j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq(n+2C_{p}logn+2\sqrt{C_{p}nlogn})\frac{C\gamma_{w}^{2}r_{n}K}{\bar{\xi}\delta_{n}})\geq 1-\frac{1}{n^{C_{p}/2}}.$ (OA.45)
Use (OA.31) and (OA.45) in (OA.34) with probability at least $1-1/n^{C_{p}/2}$ (note that we ignore the smaller order terms than the right side term in (OA.35) since they play no role in the subsequent analysis), while incorporating $\gamma_{w}^{2}$ into C,
$\max_{1\leq j\leq p}\|\hat{B}(Y_{-j}\hat{B})^{+}U_{-j}A_{j}^{+^{\prime}}a_{j}\|_{2}^{2}\leq\frac{CKr_{n}}{\bar{\eta}\bar{\xi}\delta_{n}}.$ (OA.46)
Combine all Steps 2a-2c results (OA.29)(OA.33)(OA.46) in (OA.19) to have, with probability at least $1-1/n^{C_{p}/2}$, absorbing 3 into the constant $C$,
$\max_{1\leq j\leq p}\|\hat{B}(Y_{-j}\hat{B})^{+}Fa_{j}\|_{2}^{2}\leq\frac{CK}{\bar{\xi}\delta_{n}}+\frac{\bar{\Psi}K}{\bar{\eta}\bar{\xi}\delta_{n}}+\frac{CKr_{n}}{\bar{\eta}\bar{\xi}\delta_{n}}.$ (OA.47)
Step 2d. A combination of (OA.18) (OA.47) provides the following upper bound for the first two right side terms on (OA.1), with probability at least $1-2/n^{C_{p}/2}$
$\max_{1\leq j\leq p}\|\hat{B}(Y_{-j}\hat{B})^{+}u_{j}\|_{2}^{2}$ $+$ $\|\hat{B}(Y_{-j}\hat{B})^{+}Fa_{j}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}[\frac{2logn}{n\bar{\eta}}+\frac{\bar{r}}{n\bar{\eta}}]$ (OA.48)
$+$ $\frac{CK}{\bar{\xi}\delta_{n}}+\frac{\bar{\Psi}K}{\bar{\eta}\bar{\xi}\delta_{n}}+\frac{CKr_{n}}{\bar{\eta}\bar{\xi}\delta_{n}}.$
Step 3. After setting $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j},\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$, $\bar{G}_{j}:=I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}$ definitions, we upper bound the last term on the right side of (OA.1) via Lemma 1 in the following way
$\|\alpha_{j}^{*}\|_{2}^{2}$ $=$ $\|\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}\|_{2}^{2}=\bar{a}_{j}^{\prime}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}$
$\leq$ $\frac{\bar{a}_{j}^{\prime}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}}{Eigmin(\Sigma_{U,-j})}=\frac{1}{Eigmin(\Sigma_{U,-j})}\bar{a}_{j}^{\prime}\bar{G}_{j}^{-1}(\bar{G}_{j}-I_{K})\bar{G}_{j}^{-1}\bar{a}_{j},$
Then
$\|\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}\|_{2}^{2}$$\leq$ $\frac{1}{Eigmin(\Sigma_{U,-j})}[\bar{a}_{j}^{\prime}\bar{G}_{j}^{-1}\bar{a}_{j}-\bar{a}_{j}^{\prime}\bar{G}_{j}^{-1}\bar{G}_{j}^{-1}\bar{a}_{j}]$ (OA.49)
$\leq$ $\frac{1}{Eigmin(\Sigma_{U,-j})}\bar{a}_{j}^{\prime}\bar{G}_{j}^{-1}\bar{a}_{j}.$
Next, $\bar{a}_{j}^{\prime}\bar{G}_{j}^{-1}\bar{a}_{j}=\bar{a}_{j}^{\prime}[I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}]^{-1}\bar{a}_{j}.$ Define $H_{j}:=\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}$, and $H_{j}$ is invertible, with rank K
$\bar{a}_{j}^{\prime}(I_{K}+H_{j})^{-1}\bar{a}_{j}$ $=$ $\bar{a}_{j}^{\prime}H_{j}^{-1/2}[I_{K}+H_{j}^{-1}]^{-1}H_{j}^{-1/2}\bar{a}_{j}$ (OA.50)
$\leq$ $\frac{\bar{a}_{j}^{\prime}H_{j}^{-1}\bar{a}_{j}}{Eigmin(I_{K}+H_{j}^{-1})}\leq\frac{\bar{a}_{j}^{\prime}H_{j}^{-1}\bar{a}_{j}}{1+Eigmin(H_{j}^{-1})}$
$\leq$ $\bar{a}_{j}^{\prime}H_{j}^{-1}\bar{a}_{j}=\bar{a}_{j}^{\prime}[\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}]^{-1}\bar{a}_{j},$
where we use Rayleigh inequality for the first inequality, and for the last equality we use $H_{j}$ definition. Next use $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j},\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$ definitions $\bar{a}_{j}^{\prime}[\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}]^{-1}\bar{a}_{j}=a_{j}^{\prime}(A_{-j}^{\prime}\Sigma_{U,-j}^{-1}A_{-j})^{-1}a_{j}.$ Then using the inequality after proof of (63) in Bunea et al. (2022) at the end of Section A.2.3 of Bunea et al. (2022)
$\bar{a}_{j}^{\prime}[\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}]^{-1}\bar{a}_{j}\leq\|\Sigma_{U,-j}\|_{2}a_{j}^{\prime}(A_{-j}^{\prime}A_{-j})^{-1}a_{j}.$ (OA.51)
Combine (OA.50)(OA.51) in (OA.49) to have, by (OA.25)
$\|\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}\|_{2}^{2}\leq\left[\frac{\|\Sigma_{U,-j}\|_{2}}{Eigmin(\Sigma_{U,-j})}\right]\left[\frac{a_{j}^{\prime}\Sigma_{f}a_{j}}{\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})}\right].$ (OA.52)
Next,
$\left[\frac{\|\Sigma_{U,-j}\|_{2}}{Eigmin(\Sigma_{U,-j})}\right]\left[\frac{a_{j}^{\prime}\Sigma_{f}a_{j}}{\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})}\right]$$\leq$ $\left[\frac{\|\Sigma_{U,-j}\|_{2}}{Eigmin(\Sigma_{U,-j})}\right]\left[\frac{CK}{\xi_{j}\|\Sigma_{U,-j}\|_{2}}\right]$ (OA.53)
$=$ $\frac{CK}{Eigmin(\Sigma_{U,-j})\xi_{j}}\leq\frac{CK}{\bar{\xi}},$
where we use $\min_{1\leq j\leq p}Eigmin(\Sigma_{U,-j})\geq c>0$ Assumption 2(i) for the last inequality as well as $\bar{\xi}$ definition, and second equality in (OA.26), with (OA.27).
Now we combine Steps 1-3 to have with probability at least $1-2/n^{C_{p}/2}$, with definitions of $\max_{1\leq j\leq p}\hat{r}_{j}:=\bar{r}$, $\min_{1\leq j\leq p}\hat{\eta}_{j}:=\bar{\eta}$, and $\max_{1\leq j\leq p}\hat{\Psi}_{j}:=\bar{\Psi}$,
$\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}^{2}\leq 2\gamma_{e}^{2}\sigma^{2}\left(\frac{2logn+\bar{r}}{n\bar{\eta}}\right)+\frac{CK}{\bar{\xi}}\left(1+\frac{1}{\delta_{n}}+\frac{\bar{\Psi}}{\bar{\eta}\delta_{n}}+\frac{r_{n}}{\bar{\eta}\delta_{n}}\right).$
(ii). Asymptotics is trivial when $r_{n}/\delta_{n}\to 1,$ and $\delta_{n}$ is nondecreasing in $n$.
Q.E.D.
Proof of Corollary 1. (i). We need to simplify three terms on the upper bound in Theorem 1 are $\bar{r},\bar{\eta},\bar{\Psi}$. Note that $\bar{r}:=\max_{1\leq j\leq p}\hat{r}_{j}$. In that respect,
$\hat{r}_{j}=rank(Y_{-j}P_{V_{1,k_{j}}})=rank(Y_{-j}V_{1,k_{j}})=k_{j},$ (OA.54)
where the first equality is definition and the second equality is by (OA.4), and the third equality is the proof of Corollary 5 in Bing et al. (2021). Then the term
$\hat{\eta}_{j}=\frac{1}{n}\sigma_{k_{j}}^{2}(Y_{-j}P_{V_{1,k_{j}}})=\hat{\lambda}_{k_{j}},$ (OA.55)
where the first equality is definition and the second equality is by the proof of Corollary 5 in Bing et al. (2021), and note that $\hat{\lambda}_{k_{j}}=\frac{1}{n}\sigma_{k_{j}}^{2}(Y_{-j})$ by definition. Then
$\hat{\Psi}_{j}=\frac{\sigma_{1}^{2}(Y_{-j}(I_{p-1}-P_{V_{1,k_{j}}}))}{n}=\hat{\lambda}_{k_{j}+1},$ (OA.56)
where the first equality is definition, and the second one is by the proof of Corollary 5 in Bing et al. (2021), and with $\hat{\lambda}_{k_{j}+1}=\frac{1}{n}\sigma_{k_{j}+1}^{2}(Y_{-j})$. We need to prove (21) to proceed further. First we define the event
$E_{3j}:=\{\sigma_{K+1}^{2}(Y_{-j})/n<C_{0}\Delta_{-j}\},$ (OA.57)
and its complement is $E_{3j}^{c}$. We start by proving
$P[\cap_{j=1}^{p}E_{3j}]\to 1.$ (OA.58)
In that respect, by using Weyl’s inequality
$\sigma_{K+1}(Y_{-j})\leq\sigma_{K+1}(FA_{-j}^{\prime})+\sigma_{1}(U_{-j})=\sigma_{1}(U_{-j}),$ (OA.59)
since $\sigma_{K+1}(FA_{-j}^{\prime})=0$ by $FA_{-j}^{\prime}$ having rank of $K$. See that $\sigma_{1}^{2}(U_{-j})/n=\|U_{-j}^{\prime}U_{-j}\|_{2}/n=\|U_{-j}U_{-j}^{\prime}\|_{2}/n$ by singular value definition. We provide a definition of $\tilde{U}_{-j}$, which is an $n\times p-1$ matrix, $U_{-j}$ also based on Assumption 4, with $\tilde{U}_{-jt}^{\prime}:1\times p-1$ $t$ th row of $\tilde{U}_{-j}$, and we stack these $n$ rows to get $\tilde{U}_{-j}$ and then form $U_{-j}$ $U_{-j}:=\tilde{U}_{-j}\Sigma_{U,-j}^{1/2}.$ Then by Lemma 22 of Bing et al. (2021) via Assumption 4, and definition (19)
$P[\frac{1}{n}\|\tilde{U}_{-j}\Sigma_{U,-j}\tilde{U}_{-j}^{\prime}\|_{2}\leq C_{0}\Delta_{-j}]\geq 1-exp(-n).$ (OA.60)
Define
$G_{1j}:=\{\frac{1}{n}\|\tilde{U}_{-j}\Sigma_{U,-j}\tilde{U}_{-j}^{\prime}\|_{2}\leq C_{0}\Delta_{-j}\}.$
Also see that by (OA.59)
$G_{1j}\subseteq E_{3j},$ (OA.61)
for all $j=1,\cdots,p$. By (OA.60)(OA.61)
$P(E_{3j})=P(\frac{1}{n}\sigma_{K+1}^{2}(Y_{-j})<C_{0}\Delta_{-j})\geq P(G_{1j})\geq 1-exp(-n).$ (OA.62)
So clearly $P(E_{3j}^{c})\leq exp(-n).$ To complete the proof of (OA.58)
$P(\cap_{j=1}^{p}E_{3j})=1-P((\cap_{j=1}^{p}E_{3j})^{c})=1-P(\cup_{j=1}^{p}E_{3j}^{c})\geq 1-\sum_{j=1}^{p}P(E_{3j}^{c})$
Next, the union bound argument
$P(\cup_{j=1}^{p}E_{3j}^{c})\leq\sum_{j=1}^{p}P(E_{3j}^{c})\leq exp(logp-n)\leq exp[(c^{\prime}-1)n],$ (OA.63)
given Assumption 3, $logp\leq c^{\prime}n$ with $c^{\prime}<1$ a positive constant upper bounded by one. So this clearly shows
$P(\cap_{j=1}^{p}E_{3j})\geq 1-exp((c^{\prime}-1)n).$ (OA.64)
Then by (18)(20)(OA.64), we have $P(\max_{1\leq j\leq p}\hat{s}_{j}<K+1)\geq 1-exp((c^{\prime}-1)n)$ hence
$P(\max_{1\leq j\leq p}\hat{s}_{j}\leq K)\geq 1-(exp(c^{\prime}-1)n),$ (OA.65)
with $\hat{s}_{j}$ being a positive integer. By (OA.54), $\hat{r}_{j}=k_{j}$, and since by $\hat{s}_{j}$ definition in (18) we have $k_{j}=\hat{s}_{j}$, so
$\hat{r}_{j}=\hat{s}_{j}.$ (OA.66)
Then using (OA.65) with (OA.66)
$P(\max_{1\leq j\leq p}\hat{r}_{j}\leq K)\geq 1-exp((c^{\prime}-1)n).$ (OA.67)
Now we consider $\hat{\Psi}_{j}$ uniformly in $j$ (in probability). The uniformity proof has two parts. First, define for each $j=1,\cdots,p$
$G_{2j}:=\{\hat{\lambda}_{\hat{s}_{j}+1}<C_{0}\Delta_{-j}\}.$ (OA.68)
Then first by $\hat{\Psi}_{j}$ definition in (OA.56), and since $\hat{s}_{j}=\hat{k}_{j}$ by $\hat{s}_{j}$ definition with (OA.68)
$P(\cap_{j=1}\{\hat{\Psi}_{j}<C_{0}\Delta_{-j}\})=P(\cap_{j=1}^{p}\{\hat{\lambda}_{k_{j}+1}<C_{0}\Delta_{-j}\})=P(\cap_{j=1}^{p}G_{2j}).$ (OA.69)
The first part of the uniformity proof is done. For the second part, note that by (20) and (OA.65) $\max_{1\leq j\leq p}s_{j}\leq K$ with probability approaching one
$\hat{\lambda}_{s_{j}+1}\geq\hat{\lambda}_{K+1}.$ (OA.70)
By (OA.57)(OA.68)(OA.70), and $\hat{\lambda}_{K+1}:=\frac{1}{n}\sigma_{K+1}^{2}(Y_{-j})$ with probability approaching one
$G_{2j}\subseteq E_{3j}.$ (OA.71)
So if we prove $P(\cap_{j=1}^{p}E_{3j})\to 1$, we will have $P(\cap_{j=1}^{p}G_{2j})\to 1$. We have by (OA.64) $P(\cap_{j=1}^{p}E_{3j})\geq 1-exp((c^{\prime}-1)n).$ That implies
$P(\cap_{j=1}^{p}G_{2j})\geq 1-exp((c^{\prime}-1)n).$ (OA.72)
So we proved uniformity across $j=1,\cdots,p$ for $\hat{\Psi}_{j}$ combining (OA.69) (OA.72). Now we consider $\hat{\eta}_{j}$ uniformly in $j$ (in probability).
$P(\cap_{j=1}^{p}\{\hat{\eta}_{j}\geq C_{0}\Delta_{-j}\})=P(\cap_{j=1}^{p}\{\hat{\lambda}_{\hat{s}_{j}}\geq C_{0}\Delta_{-j}\}),$ (OA.73)
since $\hat{s}_{j}=k_{j}$ and using $\hat{\eta}_{j}$ definition in (OA.55). Note that since we proved $\max_{1\leq j\leq p}\hat{s}_{j}\leq K$ with probability approaching one, and by (18)(20), uniformly over $j$ $\hat{\lambda}_{\hat{s}_{j}}\geq\hat{\lambda}_{K}.$ With (OA.73)(20) and the above inequality
$P(\cap_{j=1}^{p}\{\hat{\eta}_{j}\geq C_{0}\Delta_{-j}\})$$\geq$ $P(\cap_{j=1}^{p}\{\hat{\lambda}_{K}\geq C_{0}\Delta_{-j}\}$ (OA.74)
$=$ $P(\cap_{j=1}^{p}\{\hat{\lambda}_{K+1}<C_{0}\Delta_{-j}\})$
$\geq$ $1-exp((c^{\prime}-1)n),$
and the last inequality is by (OA.64)(OA.57), and $\hat{\lambda}_{K+1}:=\sigma_{K+1}^{2}(Y_{-j})/n$. Then $C_{0}\Delta_{-j}$ in (19)
$C_{0}\Delta_{-j}=c_{\Delta}[\|\Sigma_{U_{-j}}\|_{2}+tr(\Sigma_{U_{-j}})/n]\geq c_{\Delta}\|\Sigma_{U_{-j}\|_{2}}\geq\delta_{n},$ (OA.75)
since $c_{\Delta}>1$. Define $G_{3j}:=\{\hat{\eta}_{j}\geq\delta_{n}\}.$ Then by (OA.57) and above event definition, with equality in (OA.74) $\cap_{j=1}^{p}E_{3j}\subseteq\cap_{j=1}^{p}G_{3j},$ for each $j=1,\cdots,p$. So combine with (OA.74)(OA.75) we have
$P(\cap_{j=1}^{p}G_{3j})=P(\cap_{j=1}^{p}\{\hat{\eta}_{j}\geq\delta_{n}\})\geq P(\cap_{j=1}^{p}E_{3j})\geq 1-exp((c^{\prime}-1)n).$ (OA.76)
Furthermore, using (OA.69)(OA.74), uniformly in $j=1,\cdots,p,$ with probability approaching one
$\frac{\hat{\Psi}_{j}}{\hat{\eta}_{j}}\leq 1.$ (OA.77)
Our proofs here are uniform and detailed extensions of pointwise results of proofs of Corollaries 5-6 in Sections B.2.1-B.2.2 of Bing et al. (2021).
(ii). Note that asymptotic case is obtained by seeing $\delta_{n}$ is nondecreasing in $n$. Q.E.D.
Proof of Corollary 2. (i). We consider first $p-1<n$ scenario (in which RRE effectively becomes OLS) and again proceed by simplifying three terms, $\bar{r}$, $\bar{\eta}$, $\bar{\Psi}$, that appear in the upper bound in Theorem 1. In that respect $\hat{B}=I_{p-1}$ by definition of RRE. Then clearly $\hat{r}_{j}=rank(Y_{-j})=p-1$ since $p-1<n$ in this Corollary. Hence
$\bar{r}=\max_{1\leq j\leq p}\hat{r}_{j}<p.$ (OA.78)
Also since $I_{p-1}-P_{I_{p-1}}=0$ we have
$\bar{\Psi}=0.$ (OA.79)
This last equality shows that RRE has one less bias term in the upper bound compared to adaptive PCR estimator or compared to the general upper bound in Theorem 1. Use Corollary 11.6.7 in Bernstein (2018) (with $i=p-1,j=1$ there) or (11.6.18) of Bernstein (2018) for the first inequality below, with $\Sigma_{-j,-j}$ as defined in online Appendix B,
$\hat{\eta}_{j}$ $=$ $\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j})=\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j}\Sigma_{-j,-j}^{-1/2}\Sigma_{-j,-j}^{1/2})$ (OA.80)
$\geq$ $\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j}\Sigma_{-j,-j}^{-1/2})\sigma_{p-1}^{2}(\Sigma_{-j,-j}^{1/2})$
$=$ $[\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j}\Sigma_{-j,-j}^{-1/2})][Eigmin(\Sigma_{-j,-j})]\geq c[\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j}\Sigma_{-j,-j}^{-1/2})],$
where the last equality is by the definition of singular value. To get the last inequality in (OA.80), clearly by $\Sigma_{-j,-j},\Sigma_{U,-j}$ definitions in (OB.4), by using (5) and by Assumptions 2-3 $min_{1\leq j\leq p}Eigmin(\Sigma_{-j,-j})\geq min_{1\leq j\leq p}Eigmin(\Sigma_{U,-j})\geq c.$ So we obtain
$\min_{1\leq j\leq p}\hat{\eta}_{j}\geq c\min_{1\leq j\leq p}[\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j}\Sigma_{-j,-j}^{-1/2})].$ (OA.81)
To analyze (OA.80)(OA.81), define $D_{j}:=\{\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j}\Sigma_{-j,-j}^{-1/2})\geq C\}$.
$P(\min_{1\leq j\leq p}\frac{1}{n}\sigma_{p-1}^{2}(Y_{-j}\Sigma_{-j,-j}^{-1/2})\geq C)$ $=$ $P(\cap_{j=1}^{p}D_{j})=1-P[(\cap_{j=1}^{p}D_{j})^{c}]=1-P(\cup_{j=1}^{p}D_{j}^{c})$ (OA.82)
$\geq$ $1-\sum_{j=1}^{p}P(D_{j}^{c})\geq 1-exp(logp-C_{p}logn)$
$\geq$ $1-exp(-\frac{C_{p}}{2}logn)=1-\frac{1}{n^{C_{p}/2}},$
where to get the second inequality, with $Y_{j-1}:n\times p-1$, and $n>p-1$ here, by Theorem 5.39 of Vershynin (2012) with Corollary 10 proof in Bing et al. (2021) via Assumption 4, and applying a union bound (with $t_{1}=\sqrt{logn}$ )and by assuming $logp\leq(C_{p}/2)logn$ by Assumption 3. Next, we use (OA.80)(OA.81) and (OA.82) with $D_{j}$ definition
$P(\min_{1\leq j\leq p}\hat{\eta}_{j}\geq C)\geq 1-\frac{1}{n^{C_{p}/2}}.$ (OA.83)
Now substitute (OA.78)(OA.79)(OA.83) in the upper bound in Theorem 1 to have the result.
(ii). Asymptotics is derived by using $\delta_{n}$ is nondecreasing in $n$.
Q.E.D
Proof of Corollary 3. (i). There are again three terms that we consider, but we have $p-1>n$ in this RRE case. So
$\bar{r}=\max_{1\leq j\leq p}rank(Y_{-j})=n.$ (OA.84)
Then we have $\bar{\Psi}=0$ as in Corollary 2 due to $\hat{B}=I_{p-1}$. The last term is $\hat{\eta}_{j}=\frac{1}{n}\sigma_{n}^{2}(Y_{-j})$, which is the $n$ th largest singular value of $Y_{-j}:n\times p-1$ matrix. Also we assume $tr(\Sigma_{U,-j})/\|\Sigma_{U,-j}\|_{2}>Cn$, and $\tilde{U}_{-j}$ has independent entries. Use Lemma OB.2
$P(\hat{\eta}_{j}\geq tr(\Sigma_{U_{-j}}/n))\geq 1-4exp(-C_{p}n).$ (OA.85)
Also we have by assumption $tr(\Sigma_{U_{-j}}/n)>C\|\Sigma_{U_{-j}}\|_{2}\geq C\delta_{n},$ by Assumption 2(ii). Define $G_{4j}:=\{\hat{\eta}_{j}\geq tr(\Sigma_{U,-j}/n)\},\quad G_{5j}:=\{\hat{\eta}_{j}\geq C\delta_{n}\}.$ Note that for each $j=1,\cdots,p$
$G_{4j}\subseteq G_{5j},\mbox{hence}\quad\cap_{j=1}^{p}G_{4j}\subseteq\cap_{j=1}^{p}G_{5j}.$ (OA.86)
Note that by $G_{5j}$ definition,
$\cap_{j=1}^{p}G_{5j}=\{\min_{1\leq j\leq p}\hat{\eta}_{j}\geq C\delta_{n}\},$ (OA.87)
since the lower bounds are the same $C\delta_{n}$. Next we have the following proof
$P(\min_{1\leq j\leq p}\hat{\eta}_{j}\geq C\delta_{n})$ $=$ $P(\cap_{j=1}^{p}G_{5j})\geq P(\cap_{j=1}^{p}G_{4j})=1-P(\cup_{j=1}^{p}\{\hat{\eta}_{j}<tr(\Sigma_{U,-j})/n\})$ (OA.88)
$\geq$ $1-\sum_{j=1}^{p}P(\{\hat{\eta}_{j}<tr(\Sigma_{U,-j})/n\})\geq 1-exp(log4p-C_{p}n)$
$\geq$ $1-exp(-C_{p}n/2),$
where the first equality is by (OA.87), and the second inequality is by (OA.86), and the third inequality is by (OA.85), and the last inequality is by Assumption $3^{*}$(i). The desired result follows by incorporating these three terms in the upper bound of Theorem 1.
(ii). Asymptotics follows from (i).
Q.E.D.
Proof of Theorem 3. First of all
$\max_{1\leq j\leq p}\|\tilde{\Theta}_{j}^{\prime}-\Theta_{j}^{\prime}\|_{2}\leq\max_{1\leq j\leq p}\left|\frac{1}{\tilde{\tau}_{j}^{2}}-\frac{1}{\tau_{j}^{2}}\right|+\max_{1\leq j\leq p}\|\frac{\tilde{\alpha}_{j}}{\tilde{\tau}_{j}^{2}}-\frac{\alpha_{j}^{*}}{\tau_{j}^{2}}\|_{2}.$
Then add and subtract $\alpha_{j}^{*}/\tilde{\tau}_{j}^{2}$ to the right hand side of the immediately above inequality, and via triangle inequality and arranging
$\max_{1\leq j\leq p}\|\tilde{\Theta}_{j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$$\leq$ $\max_{1\leq j\leq p}\left|\frac{1}{\tilde{\tau}_{j}^{2}}-\frac{1}{\tau_{j}^{2}}\right|+\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}\max_{1\leq j\leq p}[\frac{1}{\tilde{\tau}_{j}^{2}}]$ (OA.89)
$+$ $\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}\max_{1\leq j\leq p}\left|\frac{1}{\tilde{\tau}_{j}^{2}}-\frac{1}{\tau_{j}^{2}}\right|.$
Note that since $\min_{1\leq j\leq p}\tilde{\tau}_{j}^{2}\geq\min_{1\leq j\leq p}\tau_{j}^{2}-\max_{1\leq j\leq p}|\tilde{\tau}_{j}^{2}-\tau_{j}^{2}|$
$\max_{1\leq j\leq p}\left(\frac{1}{\tilde{\tau}_{j}^{2}}\right)\leq\frac{1}{\min_{1\leq j\leq p}\tau_{j}^{2}-\max_{1\leq j\leq p}|\tilde{\tau}_{j}^{2}-\tau_{j}^{2}|}\leq\frac{1}{c-o_{p}(1)},$
by Theorem 2(ii)(iii) under Assumptions 1, 5. Then by (OA.53) and Assumption 5(iii)
$\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}=O(\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})=o(1).$ (OA.90)
So the third term on right side of (OA.89) converges to zero faster in probability than the first right side term in (OA.89) given Theorem 2. Then Remark 4 of Theorem 2 shows that the first term on the right side of (OA.89) converges to zero slower in probability than the second term on the same right side. So the main rate comes from the first term, which is the main diagonal term rate as shown in Theorem 2(i). Q.E.D
We now provide three corollaries for specific subcases of our general estimator. Note that we provide the specific case of adaptive PCR estimator of $\tau_{j}^{2}$, which is defined as $\hat{\tau}_{j,PCR}^{2}$. In that respect the rates $r_{w1},r_{w2}$ simplify considerably. In adaptive PCR case these rates are with wpa1
$r_{w1,PCR}=\max\left(\sqrt{\frac{logn+K}{n\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}}\right),$ (OA.91)
and
$r_{w2,PCR}=\max\left(\sqrt{\frac{logn+K}{n\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}}}\right).$ (OA.92)
See that rates (OA.91)(OA.92) can be obtained by seeing $\bar{r}\leq K,\bar{\Psi}/\bar{\eta}\leq 1,\bar{\eta}\geq\delta_{n}$ wpa1 as shown in Corollary 1. Corollary OA.7 below follows through replacing $r_{w1},r_{w2}$ in Theorem 2 with $r_{w1,PCR},r_{w2,PCR}$ respectively.
Corollary OA.7.
(i). Under Assumptions 1-5(i) and using $r_{w1,PCR},r_{w2,PCR}$ instead of $r_{w1},r_{w2}$ respectively, and $r_{n}/\delta_{n}\to 1$,
$\max_{1\leq j\leq p}|\hat{\tau}_{PCR,j}^{2}-\tau_{j}^{2}|$ $=$ $O_{p}\left(K(\sqrt{plogp/n})r_{w1,PCR}\right)$
$+$ $O_{p}\left(d_{2n}r_{w2,PCR}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\max(\sqrt{logp},\sqrt{logn})}{n^{1/2}})\right).$
(ii). Under Assumptions 1-5, $r_{n}/\delta_{n}\to 1$,
$\max_{1\leq j\leq p}|\hat{\tau}_{PCR,j}^{2}-\tau_{j}^{2}|=o_{p}(1).$
(iii). Under Assumptions 1-5, $r_{n}/\delta_{n}\to 1$,
$\min_{1\leq j\leq p}\tau_{j}^{2}\geq c-o_{p}(1).$
Here we consider the second subcase $p-1<n$ scenario of RRE estimator of $\tau_{j}^{2}$. We define it as $\hat{\tau}_{RRE,j}^{2}$. We define the following rate which will be used instead of $r_{w1},r_{w2}$, wpa1, which can be deduced from Corollary 2 proof,
$r_{RRE-l}:=\max(\sqrt{\frac{logn+p}{n}},\sqrt{\frac{K}{\bar{\xi}}}).$ (OA.93)
The proof is a simple application of Theorem 2.
Corollary OA.8.
(i).Under Assumptions 1-5(i) and using $r_{RRE-l}$ instead of both $r_{w1},r_{w2}$, setting $p/n=a_{2n}\to 0$, and $r_{n}/\delta_{n}\to 1$
$\max_{1\leq j\leq p}|\hat{\tau}_{RRE,j}^{2}-\tau_{j}^{2}|$ $=$ $O_{p}\left(K(\sqrt{a_{2n}logp})r_{RRE-l}\right)$
$+$ $O_{p}\left(d_{2n}r_{RRE-l}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\max(\sqrt{logp},\sqrt{logn})}{n^{1/2}})\right)$
(ii). Under Assumptions 1-5 with the conditions in (i)
$\max_{1\leq j\leq p}|\hat{\tau}_{RRE,j}^{2}-\tau_{j}^{2}|=o_{p}(1).$
(iii). Under Assumptions 1-5 with the conditions in (i)
$\min_{1\leq j\leq p}\tau_{j}^{2}\geq c-o_{p}(1).$
We now consider the third subcase of RRE estimator when $p-1>n$. Note that this subcase will be different than the others, and we use an estimator of $\tau_{j}^{2}$ which is different due to interpolation property of RRE in this case. First let us show the interpolation and how it affects the estimator $\tilde{\tau}_{j}^{2}$ proposed, and the need to modify. In case of $p-1>n$, RRE estimator becomes the $l_{2}$ norm interpolator. Define the minimum norm interpolator as in (1) of Bunea et al. (2022), for each $j=1,\cdots,p$ $\hat{\alpha}_{j}:=argmin\{\|\alpha_{j}\|_{2}:\|Y_{-j}\alpha_{j}-y_{j}\|_{2}=\min_{\gamma_{j}}\|Y_{-j}\gamma_{j}-y_{j}\|_{2}.\}$, Lemma 28 of Bunea et al. (2022) shows $\hat{\alpha}_{j}=Y_{-j}^{+}y_{j}=\hat{\alpha}_{RRE,j},$ where $Y_{-j}^{+}:p-1\times n$ is the Moore-Penrose inverse of $Y_{-j}$. Then when $p-1>n$ $Y_{-j}\hat{\alpha}_{j}-y_{j}=Y_{-j}Y_{-j}^{+}y_{j}-y_{j}=y_{j}-y_{j}=0,$ with $Y_{-j}Y_{-j}^{+}=I_{n}$ by (136) of Bunea et al. (2022). This is interpolation, and we cannot use the formula for $\tilde{\tau}_{j}^{2}$ since with interpolation it can be shown, $\tilde{\tau}_{j}^{2}=0$ in RRE in (12), with $p-1>n$ case. So we propose the following estimator, which can be considered a subcase of general $\tilde{\tau}_{j}^{2}$. Define
$\hat{\tau}_{INT,j}^{2}:=\frac{y_{j}^{\prime}y_{j}}{n}.$ (OA.94)
Corollary OA.9.
Under Assumptions 1-2, Assumptions $3^{*}(i)$, 3(ii), 4
(i).
$\max_{1\leq j\leq p}|\hat{\tau}_{INT,j}^{2}-\tau_{j}^{2}|=O_{p}\left(\frac{\sqrt{K}d_{2n}}{\sqrt{\bar{\xi}}}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\sqrt{logp}}{n^{1/2}})\right).$
(ii). With added to (i), Assumptions 5(ii) with $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$
$\max_{1\leq j\leq p}|\hat{\tau}_{INT,j}^{2}-\tau_{j}^{2}|=o_{p}(1).$
(iii). Using Assumptions in (ii)
$\min_{1\leq j\leq p}\tau_{j}^{2}\geq c-o_{p}(1).$
Proof of Corollary OA.9. (i). By (OB.11)(OA.94) and triangle inequality
$\max_{1\leq j\leq p}|\hat{\tau}_{INT,j}^{2}-\tau_{j}^{2}|$$\leq$ $\max_{1\leq j\leq p}|\frac{1}{n}\sum_{t=1}^{n}[y_{jt}^{2}-E(y_{jt}-\mu_{j})^{2}]|$
$+$ $\max_{1\leq j\leq p}\|E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}\|_{2}\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}.$
Next, Step 1 in the proof of Theorem 2 with (A.16), and (A.28) provides the result.
(ii). Given Assumptions in part (i) with $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$ provides the consistency.
(iii). This is the same as in the proof of Theorem 2 (iii). Q.E.D.
ONLINE APPENDIX B
In this part, we provide a precision matrix formula. Different versions with time series, and for second moments of the regressor matrix can also be found in Callot et al. (2021), Caner and Kock (2018). Its relation to a hidden factor model context is established here for the first time in the literature and in detail. We follow matrix algebra and establish a formula for each row of the precision matrix. This relation will be essential to build up our estimator. Define all assets in the portfolio as $y_{t}=(y_{jt},Y_{-jt}^{\prime})^{\prime}:p\times 1$. In case of $j=3$, this represents $y_{3t}$ in the third cell in the vector $y_{t}$ and $Y_{-3t}$ represents elements $1,2,4,\cdots,p$ in $y_{t}$.
Proof of Lemma 1. We will provide formula for $j$ th row and then stack all rows to get an expression. Before that define $\Sigma_{j,j}:=E[(y_{jt}-\mu_{j})^{2}]$ and $\Sigma_{j,-j}:=E[(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}]:1\times p-1$ and represents $j$ th row of $\Sigma$ with $j$ th element missing, and by symmetricity of $\Sigma$, $\Sigma_{j,-j}=(\Sigma_{-j,j})^{\prime}$. Specifically, define $\Sigma_{-j,j}:=E[(Y_{-jt}-\mu_{-j})(y_{jt}-\mu_{j})]:p-1\times 1$, as the $j$ th column of $\Sigma$ where $j$ th element is missing. In the same way we define $\Sigma_{-j,-j}:=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]:p-1\times p-1$ where this matrix is $j$ the row and column deleted from $\Sigma$. By Section 2.1 of Yuan (2010), the main diagonal elements are given by the scalar
$\Theta_{j,j}:=(\Sigma_{j,j}-\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j})^{-1}.$ (OB.1)
and the $j$ th row without the $j$ th element is:
$\Theta_{j,-j}:=-\Theta_{j,j}\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}.$ (OB.2)
Now we relate these quantities to a linear regression context. Define the following minimizer $\alpha_{j}^{*}:=argmin_{\alpha_{j}\in R^{p-1}}E[(y_{jt}-\mu_{j})-(Y_{-jt}-\mu_{-j})^{\prime}\alpha_{j}]^{2},$ with given zero mean $(y_{jt}-\mu_{j}),(Y_{-jt}-\mu_{-j})$ and finite second moments for $y_{jt},Y_{-jt}$, Lemma 27 in Bunea et al. (2022) shows that
$\alpha_{j}^{*}=\Sigma_{-j,-j}^{+}\Sigma_{-j,j},$ (OB.3)
where $\Sigma_{-j,-j}^{+}$ is the Moore-Penrose inverse of $\Sigma_{-j,-j}$. We will simplify that expression. First,
$\Sigma_{-j,-j}=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]=A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j},$ (OB.4)
by (3) and Assumption 1(iv). Then
$\min_{1\leq j\leq p}Eigmin(\Sigma_{-j,-j})$ $=$ $\min_{1\leq j\leq p}Eigmin(A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j})$
$\geq$ $min_{1\leq j\leq p}Eigmin(\Sigma_{U,-j})\geq c>0,$
by Assumption 2-3. So (OB.3) simplifies to
$\alpha_{j}^{*}=\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}.$ (OB.5)
Define $\eta_{jt}:=(y_{jt}-\mu_{j})-(Y_{-jt}-\mu_{-j})^{\prime}\alpha_{j}^{*}$ for $j=1,\cdots,p$. Now consider
$E[(Y_{-jt}-\mu_{-j})\eta_{jt}]$ $=$ $E[(Y_{-jt}-\mu_{-j})(y_{jt}-\mu_{j})]-E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]\alpha_{j}^{*}$ (OB.6)
$=$ $\Sigma_{-j,j}-\Sigma_{-j,-j}\alpha_{j}^{*}=\Sigma_{-j,j}-\Sigma_{-j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}=0,$
where the first equality uses $\eta_{jt}$ definition, and the second equality uses definitions of moment matrices, and the third one uses (OB.5). Given (OB.6) we can write the following infeasible linear regression model,
$y_{jt}-\mu_{j}=(Y_{-jt}-\mu_{-j})^{\prime}\alpha_{j}^{*}+\eta_{jt}.$ (OB.7)
Now we relate (OB.2) using (OB.5). Then
$\Theta_{j,-j}=-\Theta_{j,j}(\alpha_{j}^{*})^{\prime},$ (OB.8)
and then define $\tau_{j}^{2}:=E\eta_{jt}^{2}$. See that by (OB.6)(OB.7)
$E[y_{jt}-\mu_{j}]^{2}=(\alpha_{j}^{*})^{\prime}E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]\alpha_{j}^{*}+E[\eta_{jt}^{2}],$
which implies by definitions $\Sigma_{j,j}:=E[y_{jt}-\mu_{j}]^{2},\Sigma_{-j,-j}:=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]$
$\tau_{j}^{2}$ $=$ $\Sigma_{j,j}-(\alpha_{j}^{*})^{\prime}\Sigma_{-j,-j}\alpha_{j}^{*}=\Sigma_{j,j}-\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}$ (OB.9)
$=$ $1/\Theta_{j,j},$ (OB.10)
where we use (OB.5) for the second equality and (OB.1) for the third one. Clearly, also by (OB.5)(OB.9)
$\tau_{j}^{2}=\Sigma_{j,j}-\Sigma_{j,-j}\alpha_{j}^{*}.$ (OB.11)
We can form the $j$ th row of precision matrix by using $j$ the main diagonal term, via (OB.10)(OB.11)
$\Theta_{j,j}=\frac{1}{\Sigma_{j,j}-\Sigma_{j,-j}\alpha_{j}^{*}},$ (OB.12)
and the rest of that row, via (OB.8)(OB.10)
$\Theta_{j,-j}=\frac{-(\alpha_{j}^{*})^{\prime}}{\tau_{j}^{2}}.$ (OB.13)
We start simplifying $\alpha_{j}^{*}$ formula in case of hidden factors as in (2)(3). The proof for (i) here is also a detailed version of the proof of Lemma 7 and (57) of Bunea et al. (2022). This is repeated to show that our scenario fits into their framework. Note that $\Sigma_{f}:=E(f_{t}-Ef_{t})(f_{t}-Ef_{t})^{\prime}$. In that respect in (OB.5), by(2)-(3)
$\Sigma_{-j,j}$ $=$ $E[(Y_{-jt}-\mu_{-j})(y_{jt}-\mu_{j})]=E\{[A_{-j}(f_{t}-Ef_{t})+U_{-jt}][(f_{t}-Ef_{t})^{\prime}a_{j}+u_{jt}]\}$ (OB.14)
$=$ $A_{-j}\Sigma_{f}a_{j}+A_{-j}E[(f_{t}-Ef_{t})u_{jt}]+E[U_{-jt}(f_{t}-Ef_{t})^{\prime}]a_{j}+E[U_{-jt}u_{jt}]$
$=$ $A_{-j}\Sigma_{f}a_{j}=\bar{A}_{-j}\bar{a}_{j},$
by Assumption 1(iv). Then the same analysis in (OB.14) provides with $\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$,
$\Sigma_{-j,-j}=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]=A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j}=\bar{A}_{-j}\bar{A}_{-j}^{{}^{\prime}}+\Sigma_{U,-j}.$ (OB.15)
Using Sherman-Morrison-Woodbury formula in p.19 of Horn and Johnson (2013)
$\Sigma_{-j,-j}^{-1}$ $=$ $(A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j})^{-1}$ (OB.16)
$=$ $\Sigma_{U,-j}^{-1}-\Sigma_{U,-j}^{-1}A_{-j}(\Sigma_{f}^{-1}+A_{-j}^{\prime}\Sigma_{U,-j}^{-1}A_{-j})^{-1}A_{-j}^{\prime}\Sigma_{U,-j}^{-1}$
$=$ $\Sigma_{U,-j}^{-1}-\Sigma_{U,-j}^{-1}\bar{A}_{-j}(I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j})^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1},$
where we use Assumption 3, $\Sigma_{f}$ is positive definite. Note also that
$\bar{A}_{-j}^{+}\bar{A}_{-j}=\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}=I_{K},$ (OB.17)
since $A_{-j},\Sigma_{f}$ has full rank K by Assumption 3, so (135) of Bunea et al. (2022) provides the last result (OB.17). Define $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j}$ and (OB.5)(OB.14)
$\alpha_{j}^{*}$ $=$ $\Sigma_{-j,-j}^{-1}\bar{A}_{-j}\bar{a}_{j}=\Sigma_{-j,-j}^{-1}[\bar{A}_{-j}\{\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}\}\bar{a}_{j}]$ (OB.18)
$=$ $\Sigma_{-j,-j}^{-1}[\Sigma_{-j,-j}-\Sigma_{U,-j}](\bar{A}_{-j}^{+})^{\prime}\bar{a}_{j}=[I_{p-1}-\Sigma_{-j,-j}^{-1}\Sigma_{U,-j}](\bar{A}_{-j}^{+})^{\prime}\bar{a}_{j},$
where we use (OB.17) for the second equality and for the third equality (OB.15) and $\bar{A}_{-j}$ definition. By using $\bar{G}_{j}$ definition before Lemma 1 and using (OB.16) on right side of (OB.18) for $\Sigma_{-j,-j}^{-1}$
$I_{p-1}-\Sigma_{-j,-j}^{-1}\Sigma_{U,-j}$ $=$ $I_{p-1}-[\Sigma_{U,-j}^{-1}-\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}]\Sigma_{U,-j}$ (OB.19)
$=$ $\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}$
Use (OB.19) in (OB.18)
$\alpha_{j}^{*}=\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}\bar{a}_{j}=\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j},$ (OB.20)
by (OB.17) for the second equality. One of the other elements in the precision matrix is the reciprocal of the main diagonal term, $\tau_{j}^{2}$ in (OB.11). We provide a formula for that in terms of hidden factor regression model, which will be new. In that respect, by (2)(3), Assumption 1
$\Sigma_{j,j}:=E(y_{jt}-\mu_{j})^{2}=a_{j}^{\prime}\Sigma_{f}a_{j}+\sigma_{j}^{2}=\bar{a}_{j}^{\prime}\bar{a}_{j}+\sigma_{j}^{2}.$ (OB.21)
Similarly, using (2)(3), Assumption 1
$\Sigma_{j,-j}:=E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}=\bar{a}_{j}^{\prime}\bar{A}_{-j}^{\prime}.$ (OB.22)
Then combine (OB.20)(OB.21)(OB.22) in (OB.11)
$\tau_{j}^{2}=(\bar{a}_{j}^{\prime}\bar{a}_{j}+\sigma_{j}^{2})-\bar{a}_{j}^{\prime}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}.$ (OB.23)
(i)-(ii). By (OB.20) and (OB.23), and using them in (OB.10)(OB.13) we have the desired result. Q.E.D.
Here we first provide a tail inequality for quadratic form of a mean-zero subgaussian random vector. This is Lemma 21 in Bing et al. (2021).
Lemma OB.1.
Let $\zeta$ be a zero mean $d$ dimensional $\gamma_{e}$ sub-Gaussian random vector. For all symmetric positive-semi definite matrices $H$, and all $t\geq 0$,
$P\left(\zeta^{\prime}H\zeta>\gamma_{e}^{2}(\sqrt{tr(H)}+\sqrt{2\|H\|_{2}t})^{2}\right)\leq exp(-t),$
where $tr(H)$ represents trace of $H$, and $\gamma_{e}$ subgaussian random vector is defined in Assumption 4. Now, we provide a lemma for a lower bound inequality for singular value of a specific random matrix that we use in our proofs for RRE when $p-1>n$. In that way we extend the zero factor mean in Proposition 14 of Bunea et al. (2022) to a nonzero mean case below.
Lemma OB.2.
Under Assumptions $3^{*}(i)$, Assumption 3(ii), Assumption 4, with some $C>0$ large enough, $re(\Sigma_{U,-j})>Cn$, for each $j=1,\cdots,p$, and $\tilde{U}_{-j}$ has independent entries, in the case of $p-1>n$
$P\left(\sigma_{n}^{2}(Y_{-j})\geq Ctr(\Sigma_{U,-j})\right)\geq 1-4exp(-C_{p}n).$
Proof of Lemma OB.2. So first we use (89) in Bunea et al. (2022) and then equation below that on p.42 of proof of Proposition 14 of Bunea et al. (2022) but we replace Theorem 4.6.1 of Vershynin (2019) in that proof with Theorem 5.3.9 of Vershynin (2012), (set $t=\sqrt{n}$) to take into account nonzero mean of factors. The rest carries over from the proof of Proposition 14 of Bunea et al. (2022).Q.E.D.
ONLINE APPENDIX C
Simulations
By (2) and (3) and Assumption 4, for each $t=1,...,n$, we have:
$y_{t}=Af_{t}+u_{t},$ (OC.1)
where $y_{t}=(y_{jt},Y_{-jt}^{\prime})^{\prime}:p\times 1$ is the vector of excess return, $A=(a_{j},A^{\prime}_{-j}):p\times K$ is the matrix of factor loadings, and $u_{t}=(u_{jt},U_{-jt})^{\prime}:p\times 1$ vector of errors. Let hidden-factors be normally distributed and $f_{t}\sim\mathcal{N}(\mu_{f},\Sigma_{f})$, where $\mu_{f}$ is the mean vector of the hidden factor, $K\times 1$ vector with each element equal to 0.5, and $\Sigma_{f}$ is the hidden-factor covariance matrix. Following Bing et al. (2021), we set the diagonal elements of $\Sigma_{f}$ as the $K-$ length sequence from 2.5 to 3 with equal increments. For example, if $K=3$, then $\Sigma_{f}=diag(2.5,2.75,3)$. The the off diagonal, $i,j$ th elements of the factor covariance matrix are defined as:
$[\Sigma_{f}]_{i,j}=(-1)^{i+j}(0.3)^{|i-j|}\min([\Sigma_{f}]_{i,i},[\Sigma_{f}]_{j,j}).$ (OC.2)
Each element of $A$ is drawn independently from the normal distribution $\mathcal{N}(0.5,1/\sqrt{K})$. The idiosyncratic error, $u_{t}$, is drawn from the normal distribution, $\mathcal{N}(0,\Sigma_{u})$.
In our design, we define $\Sigma_{u}:p\times p$ as a diagonal matrix with diagonal elements sampled from uniform distribution with $\text{Unif}(1,3)\times\log(p)$. With that we capture noisy setups in complex models and can understand the effect of noise on adaptive PCR and RRE with $p-1>n$ specifically. The results for $K=3,5$ (true number of hidden factors) are presented in Tables OC.1-OC.2, respectively. We also have the same design with true number of latent factors as $K=20$, the results were very similar to $K=5$ hence not reported. This $K=20$ Table can be obtained from authors on demand.However note that we still use $K=20$ in other tables to see the effects of fitting larger number of factors when the true DGP has few factors. In addition, to further demonstrate how our methods behave as the signal-to-noise ratio (SNR) changes, we rewrite (OC.1) in compact form:
$Y=FA+U$ (OC.3)
Similar to Bing et al. (2021), we multiply each element of $A$ by a scalar $\alpha$ chosen from the range ${0.1,...,1.5}$. Fixing $K=20,n=400$, we experiment in high- dimensional case, with $p=450$. The low dimensional case is almost identical to high dimensional case, so we report only high dimensional case. Then for each $\alpha$, we calculate the SNR and plot the precision matrix estimation error for any of the utilized models. The results are presented in Figure 1.
[…]
Metrics
We consider metric of estimation error for the precision matrix as stated in Theorem 1. The model comparisons include several approaches from the literature. We consider the nonlinear shrinkage method of Ledoit and Wolf (2017)(NL-LW), POET method from Fan et al. (2013), and the nodewise (NW) method from Callot et al. (2021) We also tried single factor-nonlinear shrinkage method of Ledoit and Wolf (2017) but the results were worse than its no factor version of NL-LW so we decided to drop that. Our new method introduces several variations: RRE, PCR-3F, PCR-5F, PCR-20F and PCR-Adaptive. These denote the use of a RRE estimator, Principal component with fixed 3, 5, or 20 hidden factors, and an adaptive method to select the number of hidden factors, respectively. We do not include in the simulations, Caner et al. (2023) residual nodewise since it only uses observed factors. We also have done some very limited simulations with 3 factors with residual nodewise via the design here, treating unknown factors as known, and the results are similar to nodewise regression, hence omitted.
Results
Table OC.1- OC.2 shows the simulation results. The values in each cell show the average absolute estimation error for estimating the precision matrix. Each two-column block in the tables shows the result for a different sample size. The first column present results for $p=n/2$ (low-dimensional case), while the second corresponds to $p=3n/2$ (high-dimensional case). We ran 100 iterations in each simulation set up.
First, the tables show that our method achieves consistency as stated in Corollary 4 for adaptive PCR. Analyzing Table OC.1 with $K=3$, we observe that for $p=n/2$, with $n=100$, our Adaptive-PCR has an estimation error at 0.0583. This error declines to 0.0251 when $p=n/2$ and $n=400$. This indicates that as both $n$ and $p$ increase jointly, the estimation error decreases. Note that our RRE estimator in low dimensions (OLS) is consistent only when $p/n\to 0$ as shown in Corollary 5. However, it sill does well with $p=n/2$ in our Tables OC.1- OC.2. To give an example, from Table OC.1, the error is 0.4583 at $n=100$, and declines to 0.3369 at $n=400$. We also experimented with $p/n\to 0$ in our setup. As expected RRE error terms decline to zero fast, and these results are available from authors on demand.
Next, we consider which method achieves the smallest estimation errors. Our Adaptive-PCR dominates all other techniques except fixed PCR-K. In additional, our Adaptive-PCR effectively identifies the correct number of hidden factors in almost all of the scenarios. For instance, at $p=3n/2,n=400$ number of factor $K=5$ in Table OC.2, the estimation errors of both PCR-5F and Adaptive-PCR are the same at 0.0302, which also is the smallest estimation error among all methods.
Next, the number of hidden factors used in our PCR methods—whether underfitting or overfitting— impacts the estimation error. When the correct number of hidden factors is $K=3$, with $n=400$ and $p=3n/2$, the estimation error of PCR-3F is $0.0302$ in Table OC.1. However, overfitting the model by using $K=5,\,\text{or }20$ hidden factors increases the estimation error; for example, using PCR-20, the estimation error increases to $0.0401$. Underfitting, becomes problematic with $n=400$ in Table OC.2. In high dimensional scenario, PCR-5F has estimation error of 0.0302, whereas PCR-3F has error of 0.0524. As a side note to all of this, in RRE estimator, we see that estimation error for the low dimensional case is always higher than the double of high dimensional error scenario. This can be a symptom of double descent in RRE. To give an example, at Table OC.2 with $n=400$, the estimation error for RRE in low dimensional case is 0.3360, whereas in high dimensional case, RRE has error of 0.1417.
Table OC.1: Estimation Error,$K=3$
Estimator n=100 n=200 n=400
p=n/2 p=3n/2 p=n/2 p=3n/2 p=n/2 p=3n/2
NW 0.0752 0.1016 0.0617 0.0843 0.0521 0.0669
POET 47.0373 76.4717 59.6223 101.4291 84.7820 150.7152
Ledoit-Wolf 40.4218 69.1745 55.2778 97.6674 82.5374 150.9011
RRE 0.4583 0.2041 0.3858 0.1763 0.3369 0.1562
PCR-3F 0.0583 0.0738 0.0385 0.0464 0.0251 0.0302
PCR-5F 0.0641 0.0803 0.0408 0.0489 0.0263 0.0313
PCR-20F 0.1276 0.1463 0.0637 0.0712 0.0361 0.0401
PCR-Adaptive 0.0583 0.0738 0.0385 0.0464 0.0251 0.0302
Table OC.2: Estimation Error, $K=5$
Estimator n=100 n=200 n=400
p=n/2 p=3n/2 p=n/2 p=3n/2 p=n/2 p=3n/2
NW 0.0778 0.1042 0.0683 0.0888 0.0578 0.0728
POET 44.6537 72.7433 58.4469 98.6137 78.9982 135.5738
Ledoit-Wolf 38.8199 66.0089 54.5050 94.8428 76.9331 135.0054
RRE 0.4307 0.1752 0.3840 0.1577 0.3360 0.1417
PCR-3F 0.0564 0.0686 0.0488 0.0593 0.0452 0.0524
PCR-5F 0.0580 0.0732 0.0404 0.0503 0.0259 0.0302
PCR-20F 0.1193 0.1397 0.0631 0.0725 0.0355 0.0395
PCR-Adaptive 0.0559 0.0732 0.0403 0.0503 0.0259 0.0302
To examine the performance of our methods as the signal-to-noise ratio (SNR) changes, we analyze Figure 1 which shows the estimation error of the precision matrix in high-dimensional cases. We include the performance of RRE in blue and PCR-20F in red. From the figures, we observe that the performance of RRE and PCR improves as the SNR increases from 0 to 1. As the SNR increases beyond 1, the performance of PCR-20F continues to improve, while the performance of RRE stabilizes. Our PCR-20R consistently outperforms Inter-RRE across the entire SNR range for both low- and high-dimensional cases.
Figure 2: SNR and Estimation Errors: High Dimensions
Extra Empirics
Here we provide two extra tables. One for asset size 100, and the other one is for 325.
Table OC.3: Monthly Portfolio Performance of 100 Stocks, $n_{I}=240$, $n-n_{I}=108$
Estimator Maximum Sharpe Ratio
Return SD Sharpe Turnover
without transaction cost
RNW-GIC-SF 0.0072 0.0460 0.1577 0.2934
RNW-GIC-3F 0.0089 0.0463 0.1912 0.3107
RNW-GIC-5F 0.0101 0.0491 0.2055 0.3604
NW-GIC 0.0097 0.0415 0.2338 0.1554
POET 0.0080 0.0462 0.1742 0.2339
Ledoit-Wolf 0.0092 0.0540 0.1711 0.5635
Inter-RRE 0.0065 0.0646 0.1009 1.1027
PCR-1F 0.0089 0.0395 0.2265 0.1632
PCR-3F 0.0113 0.0389 0.2900 0.3368
PCR-5F 0.0079 0.0424 0.1862 0.6514
PCR-7F 0.0126 0.0477 0.2649 0.8913
PCR-Adaptive 0.0067 0.0429 0.1561 0.5223
with transaction cost
RNW-GIC-SF 0.0076 0.0457 0.1654 NA
RNW-GIC-3F 0.0092 0.0461 0.1988 NA
RNW-GIC-5F 0.0105 0.0486 0.2167 NA
NW-GIC 0.0092 0.0415 0.2218 NA
POET 0.0083 0.0461 0.1801 NA
Ledoit-Wolf 0.0095 0.0536 0.1778 NA
Inter-RRE 0.0062 0.0645 0.0963 NA
PCR-1F 0.0089 0.0397 0.2244 NA
PCR-3F 0.0112 0.0390 0.2877 NA
PCR-5F 0.0081 0.0418 0.1939 NA
PCR-7F 0.0126 0.0471 0.2685 NA
PCR-Adaptive 0.0060 0.0431 0.1396 NA
Table OC.4: Monthly Portfolio Performance of 325 Stocks, $n_{I}=240$, $n-n_{I}=108$
Estimator Maximum Sharpe Ratio
Return SD Sharpe Turnover
without transaction cost
RNW-GIC-SF 0.0047 0.0442 0.1068 0.2982
RNW-GIC-3F 0.0060 0.0451 0.1324 0.3625
RNW-GIC-5F 0.0062 0.0474 0.1300 0.4650
NW-GIC 0.0095 0.0427 0.2228 0.1393
POET 0.0057 0.0438 0.1306 0.2105
Ledoit-Wolf 0.0023 0.0554 0.0415 0.8955
Inter-RRE 0.0108 0.0457 0.2356 0.3610
PCR-1F 0.0072 0.0373 0.1933 0.1391
PCR-3F 0.0066 0.0400 0.1660 0.3270
PCR-5F 0.0064 0.0535 0.1191 1.2584
PCR-7F 0.0096 0.0449 0.2146 2.0488
PCR-Adaptive 0.0084 0.0424 0.1983 0.1841
with transaction cost
RNW-GIC-SF 0.0050 0.0439 0.1144 NA
RNW-GIC-3F 0.0062 0.0449 0.1385 NA
RNW-GIC-5F 0.0064 0.0470 0.1370 NA
NW-GIC 0.0090 0.0427 0.2111 NA
POET 0.0059 0.0439 0.1349 NA
Ledoit-Wolf 0.0022 0.0552 0.0404 NA
Inter-RRE 0.0101 0.0458 0.2205 NA
PCR-1F 0.0071 0.0375 0.1898 NA
PCR-3F 0.0071 0.0393 0.1808 NA
PCR-5F 0.0048 0.0537 0.0898 NA
PCR-7F 0.0083 0.0438 0.1905 NA
PCR-Adaptive 0.0079 0.0425 0.1857 NA